### v1

# Privacy-Preserving Federated Learning Scheme with Mitigating Model Poisoning Attacks: Vulnerabilities and Countermeasures

Abstract
The privacy-preserving federated learning schemes based on the setting of two honest-but-curious and non-colluding servers offer promising solutions in terms of security and efficiency. However, our investigation reveals that these schemes still suffer from privacy leakage when considering model poisoning attacks from malicious users. Specifically, we demonstrate that the privacy-preserving computation process for defending against model poisoning attacks inadvertently leaks privacy to one of the honest-but-curious servers, enabling it to access users’ gradients in plaintext. To address both privacy leakage and model poisoning attacks, we propose an enhanced privacy-preserving and Byzantine-robust federated learning (PBFL) scheme, comprising three components: (1) a two-trapdoor fully homomorphic encryption (FHE) scheme to bolster users’ privacy protection; (2) a novel secure normalization judgment method to preemptively thwart gradient poisoning; and (3) an innovative secure cosine similarity measurement method for detecting model poisoning attacks without compromising data privacy. Our scheme guarantees privacy preservation and resilience against model poisoning attacks, even in scenarios with heterogeneous, non-IID (Independently and Identically Distributed) datasets. Theoretical analyses substantiate the security and efficiency of our scheme, and extensive experiments corroborate the efficacy of our private attacks. Furthermore, the experimental results demonstrate that our scheme accelerates training speed while reducing communication overhead compared to the state-of-the-art PBFL schemes.
Index Terms:
Federated learning (FL), privacy protection, poisoning attacks, homomorphic encryption.
1 Introduction
Federated learning (FL) is a new distributed learning paradigm, which enables collaborative model training among a group of distributed users and a central aggregation server, while keeping users’ sensitive data local. In an ideal scenario free from distrustful entities, FL empowers users to retain complete control over their data, thereby facilitating the development of high-quality machine learning models. However, in practice, both servers and users may lack absolute trustworthiness, leading to various attacks [1, 2, 3, 4, 5, 6] against the FL framework. In recent years, the US government has been actively promoting research aimed at creating robust and privacy-preserving FL solutions for diverse tasks such as financial crime prevention [7] and pandemic response and forecasting [8]. Incentives, including substantial prizes totaling \$$\$800,000$, have been offered to spur innovation in this area. Over the years, extensive researches [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have underscored the importance of two critical security properties in FL:
  • Privacy protection of gradients: In FL, the central server may recover sensitive information of the users’ training datasets by analyzing the local gradients received from users [1, 2, 3]. To resist privacy leaks from gradients, privacy-preserving federated learning (PPFL) techniques [9, 10, 11, 12, 13] are developed. These techniques involve users masking their local gradients using privacy protection methods before submitting them to the server for secure aggregation.
  • Mitigation of poisoning gradients: Malicious users may be manipulated by Byzantine adversaries to send poisonous local gradients, deviating from correct model training. The adversarial objective may involve forcing the model to classify inputs into specific incorrect classes (targeted attacks) [4, 5] or inducing misclassifications without specifying a particular target class (untargeted attacks) [6]. Such model poisoning attacks can severely compromise the model’s performance. Recent works [15, 16, 17, 18] show that cosine similarity is an effective method to detect poisonous gradients.
The existing solutions for providing privacy and defending against poisoning attacks in FL focus on two opposing directions: privacy-preserving FL solutions aim to achieve data indistinguishability, whereas defenses against distorted gradients often rely on the similarity between malicious and benign gradients to filter out malicious ones. Specifically, defenses that extract valuable statistical information largely depend on the distinguishability of the data. Therefore, simultaneously achieving both security goals in FL remains challenging. Recently, some efforts have been made to simultaneously achieve both objectives. For instance, researchers propose privacy-preserving and Byzantine-robust FL (PBFL) schemes. However, these schemes often suffer from inefficiency [19, 20, 21, 22, 23, 24] or still make privacy leakages [22, 23, 24, 25, 26, 27]. Based on server aggregation settings, we classify existing PBFL schemes into three categories: the serverless/decentralized model, the single-server model, and the two-server model. Table I presents a qualitative comparison of FL representative works with different server aggregation settings.
TABLE I: Qualitative comparison of FL representative works.
Schemes
Privacy
Protection
Poisoning
Resilience
Server
Setting
Efficiency
FedSGD [28] $1$
PrivateFL [9, 10, 11, 12, 13] $0/1/2$
RobustFL [14, 15, 16, 17, 18] $0/1/2$
PBFL [22, 23, 24] 00
PBFL [19, 20, 21] $1$
PBFL [25, 26, 27] $2$
Our scheme $2$
  • Notes. In the column of “Server Setting”, $0,1,$ and $2$ represent the server-less, the single-server, and the two-server models, respectively.
In the serverless/decentralized model, multiple users collaborate in training a model without any central server. Blockchain techniques are commonly employed to resist model poisoning attacks in FL with the serverless setting (e.g., [22, 23, 24]) since it provides a distributed and transparent framework for storing and managing model updates in FL, thereby reducing the risk of model poisoning attacks. However, these schemes bring additional latency due to the consensus process and result in exponentially increasing communication costs among clients, thereby diminishing the efficiency of FL. Moreover, these schemes encounter challenges in privacy protection because of the openness and transparency of the data stored on the blockchain, and the introduction of privacy protection technology will bring higher storage costs and may be limited by low transaction throughput. FL schemes with the single-server model [19, 20, 21] involve a central server in the gradient aggregation process. These schemes typically employ privacy protection techniques based on multiple keys/parameters to ensure reliable privacy protection among users, such as local sensitivities of local differential privacy in [19] and individual keys of multi-key HE in [20, 21]. Despite offering privacy protection, these schemes are often inefficient. Furthermore, relying on a single server exposes it to potential single points of failure, leaving it vulnerable to attacks from external adversaries. These attacks could result in breaches of model privacy through data infringements, hacking, or leaks. FL schemes with the two-server model (e.g., [25, 26, 27]) distribute the trust of the central server in the single-server model across two non-colluding servers. These schemes present promising solutions by fortifying resilience against single points of failure and ensuring reliable privacy protection among users without necessitating multiple keys/parameters in their privacy-preserving techniques. Consequently, they hold the potential to enhance FL performance in terms of computational cost, communication burden, and final model accuracy. Nevertheless, current schemes still face challenges in effectively protecting data privacy during their secure defenses against model poisoning attacks.
In light of these challenges, this paper identifies weaknesses in PBFL schemes with the two-server model and proposes a solution. The main contributions of this paper are threefold:
(1) We analyze the security of the existing PBFL scheme based on the two-server setting and show that it violates privacy. Concretely, we present that the main secure computation procedure of the existing two-server-based scheme leaks significant information about the gradients to a semi-honest server. This vulnerability may allow the server to obtain clear gradients from all users, potentially compromising the confidentiality/privacy of the poisoning defense process.
(2) We propose an enhanced scheme based on the two-server model to protect data privacy while bolstering defenses against model poisoning attacks. In our scheme, we devise a Byzantine-tolerant aggregation method to fortify against model poisoning attacks. Additionally, we develop an enhanced secure normalization judgment $\mathbf{ESecJudge}$ method and an enhanced secure cosine similarity measurement $\mathbf{ESecCos}$ method to protect the privacy of the defense process while ensuring the correctness of the computations. Our scheme guarantees privacy preservation and resilience against model poisoning attacks, even in scenarios with heterogeneous datasets that are non-IID (Independently Identically Distributed). Furthermore, it exhibits superior computational and communication efficiency.
(3) Through experiments, we demonstrate the efficacy of our private attacks on the existing two-server-based scheme, the effectiveness of our defense against model poisoning attacks, and the efficiency of our scheme in terms of computational and communication costs. The experiment results also show that our scheme outperforms the state-of-the-art PBFL schemes.
The remainder of this paper is organized as follows. In Section 2, we introduce the related PBFL schemes. In Section 3, we introduce the system model, the threat model, and the design goals. In Section 4, we overview the two-server-based PBFL scheme ShieldFL and present our private attacks to ShieldFL. To defeat our attacks and achieve design goals, an enhanced scheme is proposed in Section 5. Section 6 analyzes the security and the complexity of the enhanced scheme. Finally, conclusions are given in Section 7. We give the necessary notations in Table II.
TABLE II: The notations and their semantic meanings.
Nots. Meanings Nots. Meanings
$\epsilon,\epsilon^{\prime}$ Security parameters $p_{1},p_{2}$ Two odd primes
$p,q$ Plaintext/ciphertext modulus $P$ Special modulus
$n_{f}$ Ring dimension $\mathcal{R}_{q}$ Cyclotomic ring
$\mathbb{R},\mathbb{C}$ Real/complex number field $\chi_{s},\chi_{e}$ Probability distribution
$pk,sk,$ $evk$ Public/secret/evaluation key $sk_{i}$ Secret key shares
[[]],delimited-[]delimited-[][\![$\cdot$]\!],[ [ ⋅ ] ] , $[\cdot]_{i}$ Ciphertext/partially decrypted ciphertext $W^{(t)}$ Model parameter, a flattened vector of all weight matrixes in a learning model
[…]
$co_{i},$ $cs_{i}$ Confidence/credit score of $U_{i}$ $\eta$ Learning rate of gradient descent method
$g_{i}^{(t)},$ $g_{\Sigma}^{(t)}$ Local/global gradient vector $g_{*}^{(t)}$ Poisonous gradient vector
$\alpha$ Parameter of Dirichlet distribution $Att_{ro}$ Attack ratio: the proportion of malicious users to the total number of users
2 Related Work
FL with the server-less/decentralized model. Blockchain techniques are usually employed to defend against poisoning attacks in this model since it provides a distributed and transparent framework for storing and managing model updates in FL, thus reducing the risk of such attacks. Kasyap et al. [22] propose a new Blockchain consensus mechanism based on honest users’ contributions on FL training, incentivizing positive training behaviors and thereby reducing the likelihood of model poisoning attacks. Xu et al. [23] propose recording an authenticator of the training program and global model on a Blockchain to resist model poisoning attacks. This scheme allows FL users to verify the global models they receive by querying the authenticator. Meanwhile, Mohamed et al. [24] design an incentive strategy that rewards reputation values to each participating user, records these values on a blockchain, and empowers users to choose users with high historical reputations for model aggregation. These schemes preserve the integrity of the federated model against poisoning attacks, but they introduce additional latency from the consensus process and incur exponentially increasing communication costs between the users, thereby reducing FL efficiency. Moreover, these schemes still face privacy protection challenges due to the openness and transparency of blockchain-stored data. Introducing privacy protection technology may escalate storage costs and could be hindered by low transaction throughput.
FL with single-server model. In general, existing researches on the single-server model achieve the secure computations of Byzantine attack detection and model aggregation by directly combining general privacy-enhancing techniques or infrastructures such as differential privacy (DP) [29, 19], homomorphic encryption (HE) [30, 20, 21], secure multiparty computation (MPC) [31, 32], and hardware-based trusted execution environment (TEE) [33]. In these schemes, the different users masked their local gradients using either a unique parameter/key (e.g., a sensitivity of global DP in [29] and a secret key of single-key HE in [30]) or using their individual parameters/keys (e.g., local sensitivities of local DP in [19] and individual keys of multi-key HE [20, 21]). The latter approach provides more reliable privacy protection but may suffer from high performance degradation, including increased communication/computation burdens and lossy model accuracy, when dealing with multi-parameter/key masked gradients. Besides, the single server is susceptible to single points of failure.
FL with the multi-server model. FL with the multi-server model, particularly the two-server model, gains attention in recent research on PBFL. In this model, users distribute their protected gradients across multiple non-colluding servers for secure aggregation. This approach mitigates single points of failure and, compared to single-server and server-less models, reduces the computational and communication burden on users without compromising model performance. While protocols with more than two servers can offer better efficiency, ensuring the validity of the honest-majority assumption remains challenges. Conversely, for dishonest-majority protocols, two-server counterparts bring more performance improvements than more servers (see, e.g., comparison in Table 6 of [34]). Liu et al. [25] propose PEFL scheme based on the two-server model. It employs linear HE to protect the privacy of local model updates and designs the corresponding secure similarity computation protocols between the two servers to detect poisonous gradient ciphertexts. However, these secure computation protocols inadvertently disclose data privacy to the servers [35]. Zhang et al. [27] propose LSFL scheme, which utilizes additive secret sharing to design a secret-sharing-based, secure Byzantine-robust aggregation protocol between the two servers. This protocol enables poison detection and secure aggregation. However, this scheme also suffers from privacy leakages during the implementation of secure aggregation between the two servers [36]. Ma et al. [26] propose ShieldFL, a two-server model employing two-trapdoor HE and cosine similarity for encrypting gradients and detecting model poisoning. In ShieldFL, the secret key of the encryption algorithm is divided into two secret shares and distributed to two non-colluding servers separately. These servers then collaborate to detect poisonous encrypted gradients and obtain an aggregation gradient using a masking technique. However, this scheme still violates privacy. The privacy breaches of these two-server-based PBFL schemes arise from the secure computation processes of similarity measurements between the two servers. Specifically, the computation involves securely computing the similarity between a local gradient and a baseline gradient (e.g., the median gradient of all gradients in [27] or the poisonous gradient baseline in ShieldFL), with the baseline gradient being identical across multiple computations. This identical baseline gradient makes it easy to design insecure schemes like ShieldFL and LSFL, which compromise privacy.
In this paper, we focus on PBFL schemes utilizing the two-server setting. These schemes fortify resilience against single points of failure and have the potential to improve FL performance regarding computational cost, communication burden, and final model accuracy. While Ma et al. [26] and Zhang et al. [27] have designed their PBFL schemes based on the two-server setting, their approaches unfortunately fall short in preserving privacy. Consequently, our scheme aims to rectify these privacy breaches and further improve computational and communication efficiency.
3 Problem Formulation
3.1 System Model
Fig. 1 illustrates the system model of our FL framework, comprising three types of entities: a key center KC, $n$ users $U_{i\in[1,n]}$, and two servers $S_{1}$ and $S_{2}$.
  • Key Center: KC generates and distributes public key and secret key shares to users and servers.
  • Users: In the $t$th training iteration, each user $U_{i}$ trains the local model $W^{(t)}$ with the individual dataset, encrypts its local gradient $g_{i}^{(t)}$ with the public key, and uploads the encrypted gradient $[\![g_{i}^{(t)}]\!]$ to $S_{1}$. Then, each user downloads the global gradient $g^{(t)}$ for local update.
  • Servers: $S_{1}$ and $S_{2}$ interact to detect poisoning of the users’ local gradients $\{[\![g_{i}^{(t)}]\!]\}_{i\in[1,n]}$ and perform aggregating of the local gradients.
Figure 1: The system model
3.2 Attack Model
Two types of attackers are considered, each with distinct objectives: Type I attackers aim to learn local gradients to compromise privacy-sensitive information, while Type II attackers seek to poison the model to degrade the final model performance. In the system, KC is a fully trusted entity. $U_{i\in[1,n]}$ are $n$ curious users, who may be benign or malicious. The benign users are Type I attackers who are only curious about other users’ data, while malicious users, in addition to this, may also be manipulated by Type II attackers to launch model poisoning attacks. We denote the set of malicious users as $U^{*}$. $S_{1}$ and $S_{2}$ are two honest-but-curious, non-colluding servers who honestly perform the given protocol but are curious about users’ privacy. The servers are assumed to be vigilant against Type II attackers, as they stand to benefit from an accurate and robust FL model.
Type I attackers: Both the honest-but-curious servers and the users are regarded as potential Type I attackers. A curious server may infer data privacy from local gradients submitted by the users. The users may obtain the other uses’ privacy-sensitive information by eavesdropping on the communication channel between the servers and the users.
Type II attackers: Byzantine adversaries are this type of attacker who manipulate malicious users to deviate from the correct model training by sending poisonous local gradients.
Given two normalized ciphertexts $[\![g_{a}]\!]=\{[\![x_{a_{1}}]\!],\cdots,[\![x_{a_{l}}]\!]\}$ and $[\![g_{b}]\!]=\{[\![x_{b_{1}}]\!],\cdots,[\![x_{b_{l}}]\!]\}$, $S_{1}$ interacts with $S_{2}$ to obtain the cosine similarity $\cos_{ab}$ of $g_{a}$ and $g_{b}$:
  • $\bullet$
    @$S_{1}$: On receiving $[\![g_{a}]\!]$ and $[\![g_{b}]\!]$, $S_{1}$ selects $l$ random noises $r_{i\in[1,l]}\leftarrow\mathbb{Z}_{N}^{*}$ and masks $[\![g_{a}]\!]$ and $[\![g_{b}]\!]$ as $[\![\bar{g}_{a}]\!]=\{[\![\bar{x}_{a_{1}}]\!],\cdots,[\![\bar{x}_{a_{l}}]\!]\}$ and $[\![g_{b}]\!]=\{[\![\bar{x}_{b_{1}}]\!],\cdots,[\![\bar{x}_{b_{l}}]\!]\}$, where $[\![\bar{x}_{a_{i}}]\!]=[\![x_{a_{i}}]\!]\cdot[\![r_{i}]\!]$ and $[\![\bar{x}_{b_{i}}]\!]=[\![x_{b_{i}}]\!]\cdot[\![r_{i}]\!]$. Then, $S_{1}$ partially decrypts the masked ciphertexts as $[\bar{x}_{a_{i}}]_{1}\leftarrow\mathbf{PHE.PartDec}_{sk_{1}}([\![\bar{x}_{a_{i}}]\!])$ and $[\bar{x}_{b_{i}}]_{1}\leftarrow\mathbf{PHE.PartDec}_{sk_{1}}([\![\bar{x}_{b_{i}}]\!]),s.t.,[\![\bar{x}_{a_{i}}]\!]\in[\![\bar{g}_{a}]\!],[\![\bar{x}_{b_{i}}]\!]\in[\![\bar{g}_{b}]\!]$, and sends $[\bar{g}_{a}]_{1},[\bar{g}_{b}]_{1},[\![\bar{g}_{a}]\!],[\![\bar{g}_{b}]\!]$ to $S_{2}$.
  • $\bullet$
    @$S_{2}$: Once obtaining the above ciphertexts, $S_{2}$ calls $\mathbf{PHE.PartDec}$ and $\mathbf{PHE.FullDec}$ to perform full decryption and obtain the masked gradients as $\bar{g}_{a}$ and $\bar{g}_{b}$. Then $S_{2}$ calculates the cosine similarity between $\bar{g}_{a}$ and $\bar{g}_{b}$ as $\overline{\cos}_{ab}=\bar{g}_{a}\odot\bar{g}_{b}=\sum_{i=1}^{m}\bar{x}_{a_{i}}\cdot\bar{x}_{b_{i}}$, where $\odot$ represents inner product. $S_{2}$ calls the encryption algorithm $\mathbf{PHE.Enc}$ to return $[\![\overline{\cos}_{ab}]\!]$ to $S_{1}$.
  • $\bullet$
    @$S_{1}$: $S_{1}$ removes the noise $r_{i}$ in $[\![\overline{\cos}_{ab}]\!]$ to obtain the final cosine similarity $[\![\cos_{ab}]\!]$ and sends it to $S_{2}$.
  • $\bullet$
    @$S_{2}$: $S_{2}$ computes the decryption share $[\cos_{ab}]_{2}\leftarrow\mathbf{PHE.PartDec}_{sk_{2}}([\![\cos_{ab}]\!])$ and returns it to $S_{1}$.
  • $\bullet$
    @$S_{1}$: On receiving $[\cos_{ab}]_{2}$, $S_{1}$ obtains the cosine similarity plaintext $\cos_{ab}$ by calling $\mathbf{PHE.PartDec}$ and $\mathbf{PHE.FullDec}$.
Fig. 2 Implementation procedure of $\mathbf{SecCos}([\![g_{a}]\!],[\![g_{b}]\!])$.
3.3 Design Goals
The design goals are to ensure security and performance as detailed below.
  • Confidentiality (Privacy Protection). The scheme should be resistant to Type I attackers, ensuring that users’ private datasets cannot be reconstructed by this type of attackers (including other users and servers).
  • Robustness. The scheme should be resistant to Type II attackers. It should be capable of detecting model poisoning attacks from malicious users based on users’ encrypted local models and should be able to produce a high-quality global model.
  • Efficiency. The computational and communication overhead on users and servers should be minimized.
4 Security Vulnerabilities of Existing PBFL Schemes
In this section, we briefly review the existing two-server-based PBFL scheme ShieldFL [26] and identify its security vulnerabilities. We begin by discussing the technical background of ShieldFL, followed by outlining its detailed processes. Finally, we present the privacy breaches observed in ShieldFL.
4.1 Technical Background
ShieldFL utilizes a two-trapdoor partial homomorphic encryption (PHE) scheme based on Paillier’s cryptosystem to protect privacy, which contains five algorithms:
  • $\mathbf{PHE.KeyGen}(\varepsilon)\rightarrow(pk,sk)$: On input of the security parameter $\varepsilon$, the key generation algorithm first generates two odd primes $p_{1}$ and $p_{2}$ such that $|p_{1}|=|p_{2}|=\varepsilon$ and sets $N=p_{1}\cdot p_{2}$. Then, it outputs the public key $pk=(N,1+N)$ and the secret key $sk=\text{lcm}(p_{1}-1,p_{2}-1)$.
  • $\mathbf{PHE.Enc}_{pk}(m)\rightarrow[\![m]\!]$: On input of a plaintext $m\in\mathbb{Z}_{N}$, the encryption algorithm encrypts $m$ with $pk$ and outputs the ciphertext $[\![m]\!]=(1+N)^{m}\cdot r^{N}\bmod N^{2}$, where $r\in\mathbb{Z}^{*}_{N}$.
  • $\mathbf{PHE.KeySplit}(sk)\rightarrow(sk_{1},sk_{2})$: Given the secret key $sk$, two secret key shares $sk_{1}$ and $sk_{2}$ are generated through randomly dividing $sk$ meeting $\sum_{i=1}^{2}sk_{i}\equiv 0\bmod sk$ and $\sum_{i=1}^{2}sk_{i}\equiv 1\bmod N.$
  • $\mathbf{PHE.PartDec}_{sk_{i}}([\![m]\!])\rightarrow[m]_{i}$: Given a ciphertext $[\![m]\!]$, a decryption share $[m]_{i}=[\![m]\!]^{sk_{i}}$ is output through partially decrypting $[\![m]\!]$ with the secret key share $sk_{i}$.
  • $\mathbf{PHE.FullDec}([m]_{1},[m]_{2})\rightarrow m$: Given the decryption share pair $([m]_{1},[m]_{2})$, the fully decrypted plaintext $m=\frac{(\prod_{i=1}^{2}[m]_{i}\bmod N^{2})-1}{N}\bmod N$ is output.
More details of the five algorithms can be referred to [26]. The two-trapdoor PHE scheme meets the additive homomorphism such that $[\![m_{1}]\!]+[\![m_{2}]\!]=[\![m_{1}+m_{2}]\!]$ and $[\![m]\!]^{k}=[\![km]\!]$.
4.2 ShieldFL Framework
ShieldFL consists of four processes:
  1. 1.
    Setup: In this process, KC and $S_{1}$ distribute keys and an initial model, respectively.
    • KC generates a public/secret key pair $(pk,sk)$ and calls $\mathbf{PHE.KeySplit}(sk)$ with $n+1$ times to generate $n+1$ pairs of independent secret key shares $(sk_{1},sk_{2})$, $(sk_{s_{1}},sk_{u_{1}}),\cdots,(sk_{s_{n}},sk_{u_{n}})$. Then, KC broadcasts $pk$; distributes $sk_{1}$ and $sk_{2}$ to $S_{1}$ and $S_{2}$, respectively; and distributes $\{sk_{s_{i}}\}_{i\in[1,n]}$ and $sk_{u_{i}}$ to $S_{1}$ and $U_{i}$, respectively.
    • $S_{1}$ releases the encrypted initial model $[\![W^{(0)}]\!]$ and its decryption share $[W^{(0)}]_{s_{i}}\leftarrow\mathbf{PHE.PartDec}_{sk_{s_{i}}}([\![W^{(0)}]\!])$ to $U_{i}$ for $i\in[1,n]$.
    • Each user $U_{i}$ calculates the decryption share $[W^{(0)}]_{u_{i}}$leftarrow $\mathbf{PHE.PartDec}_{sk_{u_{i}}}([\![W^{(0)}]\!])$ and then obtains the fully decrypted initial model $W^{(0)}$leftarrow $\mathbf{PHE.FullDec}$ $([W^{(0)}]_{s_{i}},[W^{(0)}]_{u_{i}})$ to update its local model.
  2. 2.
    Local Training: In each training iteration, each user trains its local model to obtain a local gradient, encrypts the local gradient, and sends the encrypted gradient to $S_{1}$. Specifically, in the $t$th iteration, for a begin user $U_{i}$, it trains the local model $W^{(t)}$ and normalizes its local gradient $g_{i}^{(t)}$, then it sends the encrypted normalized gradient $[\![g_{i}^{(t)}]\!]$ to $S_{1}$. For a malicious user $U_{i}\in U^{*}$, it launches model poisoning and sends encrypted poisonous gradients $[\![g_{i*}^{(t)}]\!]$ to $S_{1}$.
  3. 3.
    Privacy-Preserving Defense: This process mainly consists of the following three implementations.
    • Normalization judgment. $S_{1}$ determines whether it received encrypted local gradients $[\![g_{i}^{(t)}]\!]$ are normalized by implementing a normalization judgment procedure $\mathbf{SecJudge}([\![g_{i}^{(t)}]\!])$.
    • Poisonous gradient baseline finding. $S_{1}$ finds the gradient $[\![g_{*}^{(t)}]\!]$ (among the users’ local gradients) with the lowest cosine similarity as the baseline of poisonous gradients through implementing a secure cosine similarity measurement $\cos_{i\Sigma}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!])$, where $[\![g_{\Sigma}^{(t-1)}]\!]$ is the aggregated gradient of the last iteration.
    • Byzantine-tolerance aggregation. First, $S_{1}$ measures the cosine similarity between the poisonous baseline and each user’s local gradient as $\cos_{i*}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!]).$ Then, $S_{1}$ sets the confidence $co_{i}\leftarrow 1\times deg-\cos_{i*}$ for $g_{i}^{(t)}$ and calculates $co\leftarrow\sum_{i=1}^{n}co_{i},co_{i}=\lfloor\dfrac{co_{i}}{co}\cdot deg\rceil$. Finally, $S_{1}$ weights the local gradients with their respective confidences and then aggregates them to generate a global gradient $[\![g_{\Sigma}^{(t)}]\!]$, where $g_{\Sigma}^{(t)}=\sum_{i=1}^{n}co_{i}g_{i}^{(t)}$.
  4. 4.
    Model Update. $S_{1}$ sends $[\![g_{\Sigma}^{(t)}]\!]$ and $[g_{\Sigma}^{(t)}]_{s_{i}}\leftarrow\mathbf{PHE.PartDec}_{sk_{s_{i}}}([\![g_{\Sigma}^{(t)}]\!])$ to $U_{i}$. Then $U_{i}$ decrypts to obtain the aggregated gradient plaintext $g_{\Sigma}^{(t)}$ and updates its local model as $W^{(t+1)}\leftarrow W^{(t)}-\eta^{(t)}g_{\Sigma}^{(t)}$, where $\eta^{(t)}$ is the learning rate of the $t$th training iteration.
In ShieldFL, privacy leakages primarily occur during the second and third implementations, namely, the poisonous gradient baseline finding and Byzantine-tolerance aggregation, of the privacy-preserving defense process. Therefore, we next provide a detailed procedure of the secure computation involved in these two implementations, as illustrated in Fig. 2: the implementation procedure of the secure cosine similarity measurement method $\mathbf{SecCos}$. Further information on secure computations such as $\mathbf{SecJudge}$, please refer to [26]. For further information on other secure computation procedures such as $\mathbf{SecJudge}$, please refer to [26].
4.3 Privacy Leakage in ShieldFL
4.3.1 Poisonous Gradient Baseline Finding
In the procedure of poisonous gradient baseline finding, $S_{1}$ interacts with $S_{2}$ to measure the secure cosine similarity $\cos_{i\Sigma}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!])$ for $i\in[1,n]$, so as to find the baseline of poisonous gradient $[\![g_{*}^{(t)}]\!]\in\{[\![g_{i}^{(t)}]\!]\}_{i\in[1,n]}$ with the lowest cosine similarity. Specifically, in this procedure, $S_{1}$ selects $l$ random noises $\{r_{i1},\cdots,r_{il}\}$ and masks $[\![g_{i}^{(t)}]\!]=\{[\![x_{i1}]\!],\cdots,[\![x_{il}]\!]\}$ and $[\![g_{\Sigma}^{(t-1)}]\!]=\{[\![y_{1}]\!],\cdots,[\![y_{l}]\!]\}$ as $[\![\bar{g}_{i}^{(t)}]\!]$ and $[\![\bar{g}_{\Sigma,i}^{(t)}]\!]$ using Eq. (1) and Eq. (2), respectively. Then $S_{1}$ partially decrypts $[\![\bar{g}_{i}^{(t)}]\!]$ and $[\![\bar{g}_{\Sigma,i}^{(t)}]\!]$ as $[\bar{g}_{i}^{(t)}]_{1}$ and $[\bar{g}_{\Sigma,i}^{(t)}]_{1}$ and sends $[\![\bar{g}_{i}^{(t)}]\!],[\![\bar{g}_{\Sigma,i}^{(t)}]\!],[\bar{g}_{i}^{(t)}]_{1},[\bar{g}_{\Sigma,i}^{(t)}]_{1}$ to $S_{2}$.
$\small\left(\begin{array}[]{c}[\![\bar{g}_{1}^{(t)}]\!]\\ \vdots\\ {[\![\bar{g}_{i}^{(t)}]\!]}\\ \vdots\\ {[\![\bar{g}_{n}^{(t)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccc}[\![x_{11}]\!]\cdot[\![r_{11}]\!]&\cdots&[\![x_{1l}]\!]\cdot[\![r_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{i1}]\!]\cdot[\![r_{i1}]\!]}&\cdots&[\![x_{il}]\!]\cdot[\![r_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{n1}]\!]\cdot[\![r_{n1}]\!]}&\cdots&[\![x_{nl}]\!]\cdot[\![r_{nl}]\!]\\ \end{array}\right)$ (1)
$\small\left(\begin{array}[]{c}[\![\bar{g}_{\Sigma,1}^{(t-1)}]\!]\\ \vdots\\ {[\![\bar{g}_{\Sigma,i}^{(t-1)}]\!]}\\ \vdots\\ {[\![\bar{g}_{\Sigma,n}^{(t-1)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccccc}[\![y_{1}]\!]\cdot[\![r_{11}]\!]&\cdots&[\![y_{l}]\!]\cdot[\![r_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![y_{1}]\!]\cdot[\![r_{i1}]\!]}&\cdots&[\![y_{l}]\!]\cdot[\![r_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![y_{1}]\!]\cdot[\![r_{n1}]\!]}&\cdots&[\![y_{l}]\!]\cdot[\![r_{nl}]\!]\\ \end{array}\right)$ (2)
Privacy Leakage: Once receiving the ciphertexts $[\![\bar{g}_{i}^{(t)}]\!],[\![\bar{g}_{\Sigma,i}^{(t)}]\!],[\bar{g}_{i}^{(t)}]_{1},[\bar{g}_{\Sigma,i}^{(t)}]_{1}$ from $S_{1}$, the server $S_{2}$ calls $\mathbf{PHE.PartDec}$ and $\mathbf{PHE.FullDec}$ to obtain $\bar{g}_{i}^{(t)}$ and $\bar{g}_{\Sigma,i}^{(t-1)}$ as Eq. (3) and Eq. (4), respectively.
$\small\left(\begin{array}[]{c}\bar{g}_{1}^{(t)}\\ \vdots\\ {\bar{g}_{i}^{(t)}}\\ \vdots\\ {\bar{g}_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}x_{11}+r_{11}&\cdots&x_{1l}+r_{1l}\\ \vdots&\cdot&\vdots\\ {x_{i1}+r_{i1}}&\cdots&x_{il}+r_{il}\\ \vdots&\cdot&\vdots\\ {x_{n1}+r_{n1}}&\cdots&x_{nl}+r_{nl}\\ \end{array}\right)$ (3)
$\small\left(\begin{array}[]{c}\bar{g}_{\Sigma,1}^{(t-1)}\\ \vdots\\ {\bar{g}_{\Sigma,i}^{(t-1)}}\\ \vdots\\ {\bar{g}_{\Sigma,n}^{(t-1)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}y_{1}+r_{11}&\cdots&y_{l}+r_{1l}\\ \vdots&\cdot&\vdots\\ {y_{1}+r_{i1}}&\cdots&y_{l}+r_{il}\\ \vdots&\cdot&\vdots\\ {y_{1}+r_{n1}}&\cdots&y_{l}+r_{nl}\\ \end{array}\right)$ (4)
Then, $S_{2}$ can learn the difference between two arbitrary local gradients. The difference between the local gradients of the $i$th and $j$th users is calculated as
$\begin{split}d_{ij}&=g_{i}^{(t)}-g_{j}^{(t)}=\{x_{i1}-x_{j1},\cdots,x_{il}-x_{jl}\}\\ &=\left(\bar{g}_{i}^{(t)}-\bar{g}_{j}^{(t)}\right)-\left(\bar{g}_{\Sigma,i}^{(t-1)}-\bar{g}_{\Sigma,j}^{(t-1)}\right).\end{split}$ (5)
Besides, $S_{2}$ can learn the difference between the last-iteration aggregated gradient $g_{\Sigma}^{(t-1)}$ and an arbitrary local gradient $g_{i}^{(t)}$:
$\begin{split}d_{i\Sigma}&=g_{i}^{(t)}-g_{\Sigma}^{(t-1)}=\{x_{i1}-y_{1},\cdots,x_{il}-y_{l}\}\\ &=\bar{g}_{i}^{(t)}-\bar{g}_{\Sigma,i}^{(t-1)}.\end{split}$ (6)
Thus, according to Eqs. (5) and (6), $S_{2}$ knows
$\small\left(\begin{array}[]{c}g_{1}^{(t)}\\ \vdots\\ {g_{i}^{(t)}}\\ \vdots\\ {g_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{c}d_{1j}+g_{j}^{(t)}\\ \vdots\\ {d_{ij}+g_{j}^{(t)}}\\ \vdots\\ {d_{nj}+g_{j}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{c}d_{1\Sigma}+g_{\Sigma}^{(t-1)}\\ \vdots\\ {d_{i\Sigma}+g_{\Sigma}^{(t-1)}}\\ \vdots\\ {d_{n\Sigma}+g_{\Sigma}^{(t-1)}}\\ \end{array}\right),$ (7)
in which $S_{2}$ acquires a “shifted” distribution of all local gradients in plaintext (It is worth noting that $\{d_{ij}\}_{i\in[1,n],j\in[1,n]}$ and $\{d_{i\Sigma}\}_{i\in[1,n]}$ are all accessible to $S_{2}$). This compromises privacy significantly [2], as it divulges a considerable amount of information to $S_{2}$, thereby failing to fulfill the confidentiality requirement outlined in [26].
4.3.2 Byzantine-Tolerance Aggregation
$S_{1}$ interacts with $S_{2}$ to implement the secure cosine similarity measurement $\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!])$ for $i=1,\cdots,n$, so as to obtain the similarities of all local gradients related to the poisonous gradient baseline $[\![g_{*}^{(t)}]\!]$. Specifically, in this procedure, $S_{1}$ selects $l$ random noises $\{\gamma_{i1},\cdots,\gamma_{il}\}$ and masks $[\![g_{i}^{(t)}]\!]=\{[\![x_{i1}]\!],\cdots,[\![x_{il}]\!]\}$ and $[\![g_{*}^{(t)}]\!]=\{[\![z_{1}]\!],\cdots,[\![z_{l}]\!]\}$ as $[\![\tilde{g}_{i}^{(t)}]\!]$ and $[\![\tilde{g}_{*,i}^{(t)}]\!]$ using Eq. (8) and Eq. (9), respectively. Then $S_{1}$ partially decrypts $[\![\tilde{g}_{i}^{(t)}]\!]$ and $[\![\tilde{g}_{*,i}^{(t)}]\!]$ as $[\tilde{g}_{i}^{(t)}]_{1}$ and $[\tilde{g}_{*,i}^{(t)}]_{1}$ and sends $[\![\tilde{g}_{i}^{(t)}]\!],[\![\tilde{g}_{*,i}^{(t)}]\!],[\tilde{g}_{i}^{(t)}]_{1},[\tilde{g}_{*,i}^{(t)}]_{1}$ to $S_{2}$.
$\small\left(\begin{array}[]{c}[\![\tilde{g}_{1}^{(t)}]\!]\\ \vdots\\ {[\![\tilde{g}_{i}^{(t)}]\!]}\\ \vdots\\ {[\![\tilde{g}_{n}^{(t)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccc}[\![x_{11}]\!]\cdot[\![\gamma_{11}]\!]&\cdots&[\![x_{1l}]\!]\cdot[\![\gamma_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{i1}]\!]\cdot[\![\gamma_{i1}]\!]}&\cdots&[\![x_{il}]\!]\cdot[\![\gamma_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{n1}]\!]\cdot[\![\gamma_{n1}]\!]}&\cdots&[\![x_{nl}]\!]\cdot[\![\gamma_{nl}]\!]\\ \end{array}\right)$ (8)
$\small\left(\begin{array}[]{c}[\![\tilde{g}_{*,1}^{(t-1)}]\!]\\ \vdots\\ {[\![\tilde{g}_{*,i}^{(t-1)}]\!]}\\ \vdots\\ {[\![\tilde{g}_{*,n}^{(t-1)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccccc}[\![z_{1}]\!]\cdot[\![\gamma_{11}]\!]&\cdots&[\![z_{l}]\!]\cdot[\![\gamma_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![z_{1}]\!]\cdot[\![\gamma_{i1}]\!]}&\cdots&[\![z_{l}]\!]\cdot[\![\gamma_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![z_{1}]\!]\cdot[\![\gamma_{n1}]\!]}&\cdots&[\![z_{l}]\!]\cdot[\![\gamma_{nl}]\!]\\ \end{array}\right)$ (9)
Privacy Leakage: Upon obtaining the ciphertexts $[\![\tilde{g}_{i}^{(t)}]\!],[\![\tilde{g}_{*,i}^{(t)}]\!],[\tilde{g}_{i}^{(t)}]_{1},[\tilde{g}_{*,i}^{(t)}]_{1}$ from $S_{1}$, the server $S_{2}$ calls $\mathbf{PHE.PartDec}$ and $\mathbf{PHE.FullDec}$ to obtain $\tilde{g}_{i}^{(t)}$ and $\tilde{g}_{*,i}^{(t)}$ as Eq. (10) and Eq. (11), respectively.
$\small\left(\begin{array}[]{c}\tilde{g}_{1}^{(t)}\\ \vdots\\ {\tilde{g}_{i}^{(t)}}\\ \vdots\\ {\tilde{g}_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}x_{11}+\gamma_{11}&\cdots&x_{1l}+\gamma_{1l}\\ \vdots&\cdot&\vdots\\ {x_{i1}+\gamma_{i1}}&\cdots&x_{il}+\gamma_{il}\\ \vdots&\cdot&\vdots\\ {x_{n1}+\gamma_{n1}}&\cdots&x_{nl}+\gamma_{nl}\\ \end{array}\right)$ (10)
$\vspace*{-0.5\baselineskip}\small\left(\begin{array}[]{c}\tilde{g}_{*,1}^{(t)}\\ \vdots\\ {\tilde{g}_{*,i}^{(t)}}\\ \vdots\\ {\tilde{g}_{*,n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}z_{1}+\gamma_{11}&\cdots&z_{l}+\gamma_{1l}\\ \vdots&\cdot&\vdots\\ {z_{1}+\gamma_{i1}}&\cdots&z_{l}+\gamma_{il}\\ \vdots&\cdot&\vdots\\ {z_{1}+\gamma_{n1}}&\cdots&z_{l}+\gamma_{nl}\\ \end{array}\right)$ (11)
Since $\tilde{g}_{*}^{(t)}$ is the gradient of a certain user, i.e., $\tilde{g}_{*}^{(t)}\in\{\tilde{g}_{1}^{(t)},\cdots,\tilde{g}_{n}^{(t)}\}$, $S_{2}$ can learn which user’s gradient is the poisonous gradient baseline by comparing $\tilde{g}_{*,i}^{(t)}$ and $\tilde{g}_{i}^{(t)}$. That is, if $\tilde{g}_{*,k}^{(t)}=\tilde{g}_{k}^{(t)}$, then $S_{2}$ can determine that the gradient of the $k$th user is the poisonous gradient baseline which has the lowest cosine similarity in the $t$th iteration of the model. This, to some extent, leaks the $k$th user’s personal information. Besides, similar to the analyses in the poisonous gradient baseline finding, in Byzantine-tolerance aggregation, $S_{2}$ can also learn the difference between arbitrary two local gradients from Eq. (10) and Eq. (11). Therefore, $S_{2}$ knows the differences between the gradient of an arbitrary user $i$ and the poisonous gradient baseline, which is
$\begin{split}d_{i*}&=g_{i}^{(t)}-g_{*}^{(t)}=\{x_{i1}-z_{1},\cdots,x_{il}-z_{l}\}\\ &=\left(\tilde{g}_{i}^{(t)}-\tilde{g}_{l}^{(t)}\right)-\left(\tilde{g}_{*,i}^{(t)}-\tilde{g}_{*,l}^{(t)}\right).\end{split}$ (12)
Thus, $S_{2}$ knows a “shifted” distribution of all user gradients related to the poisonous gradient baseline through Eq. (13).
$\small\left(\begin{array}[]{c}g_{1}^{(t)}\\ \vdots\\ {g_{i}^{(t)}}\\ \vdots\\ {g_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{c}d_{1*}+g_{*}^{(t)}\\ \vdots\\ {d_{i*}+g_{*}^{(t)}}\\ \vdots\\ {d_{n*}+g_{*}^{(t)}}\\ \end{array}\right).$ (13)
In addition, $S_{2}$ can learn the difference between the poisonous gradient baseline and the last-iteration aggregated gradient:
$\begin{split}d_{*\Sigma}&=g_{*}^{(t)}-g_{\Sigma}^{(t-1)}=\{z_{1}-y_{1},\cdots,z_{l}-y_{l}\}\\ &=\left(\tilde{g}_{*,i}^{(t)}-\bar{g}_{\Sigma,i}^{(t-1)}\right)-\left(\tilde{g}_{i}^{(t)}-\bar{g}_{i}^{(t)}\right).\end{split}$ (14)
Thus, the correlation between the poisonous gradient baseline $g_{*}^{(t)}$ and the last-iteration aggregated gradient $g_{\Sigma}^{(t-1)}$ is leaked to $S_{2}$, which is $g_{*}^{(t)}=g_{\Sigma}^{(t-1)}+d_{*\Sigma}$ according to Eq. (14).
4.3.3 Gradient Leakage in Clear
Based on the privacy breaches mentioned above, the local gradients can be directly leaked in plaintext, thus breaking the privacy of the whole ShieldFL scheme. Specifically, once $S_{2}$ knows one of the gradients in $\{g_{*}^{(t)},g_{\Sigma}^{(t-1)},\{g_{i}^{(t)}\}_{i\in[1,n]}\}$, it can directly obtain all other users’ local gradients in clear using Eq. (7) or Eq. (13). We present that $S_{2}$ can obtain one of these gradients using the following methods.
  1. a)
    Since $\{g_{*}^{(t)},g_{\Sigma}^{(t-1)},\{g_{i}^{(t)}\}_{i\in[1,n]}\}$ are all normalized gradients, and the difference between any two of these gradients is a known value, it is sufficient to $S_{2}$ to guess a gradient in $\{g_{*}^{(t)},g_{\Sigma}^{(t-1)},\{g_{i}^{(t)}\}_{i\in[1,n]}\}$. For example, $S_{2}$ knows that
    $|\!|g_{j}^{(t)}|\!|={{x}_{j1}^{2}}+\cdots+{x}_{jl}^{2}=1.$ (15)
    According to Eq. (7), $S_{2}$ knows that
    $\tiny\left(\begin{array}[]{c}|\!|g_{1}^{(t)}|\!|\\ \vdots\\ {|\!|g_{i}^{(t)}|\!|}\\ \vdots\\ {|\!|g_{n}^{(t)}|\!|}\\ \end{array}\right)=\left(\begin{array}[]{c}|\!|d_{1j}+g_{j}^{(t)}|\!|\\ \vdots\\ {|\!|d_{ij}+g_{j}^{(t)}|\!|}\\ \vdots\\ {|\!|d_{nj}+g_{j}^{(t)}|\!|}\\ \end{array}\right)=\left(\begin{array}[]{c}\sum_{k=1}^{l}(d_{1j}(k)+x_{jk})^{2}\\ \vdots\\ {\sum_{k=1}^{l}(d_{ij}(k)+x_{jk})^{2}}\\ \vdots\\ {\sum_{k=1}^{l}(d_{nj}(k)+x_{jk})^{2}}\\ \end{array}\right)=\left(\begin{array}[]{c}1\\ \vdots\\ 1\\ \vdots\\ 1\\ \end{array}\right),$ (16)
    $\tiny\begin{split}&\left(\begin{array}[]{c}|\!|g_{\Sigma}^{(t-1)}|\!|\\ \vdots\\ {|\!|g_{\Sigma}^{(t-1)}|\!|}\\ \vdots\\ {|\!|g_{\Sigma}^{(t-1)}|\!|}\\ \end{array}\right)=\left(\begin{array}[]{c}|\!|d_{1j}+g_{j}^{(t)}-d_{1\Sigma}|\!|\\ \vdots\\ {|\!|d_{ij}+g_{j}^{(t)}-d_{i\Sigma}|\!|}\\ \vdots\\ {|\!|d_{nj}+g_{j}^{(t)}-d_{n\Sigma}|\!|}\\ \end{array}\right)\\ =&\left(\begin{array}[]{c}\sum_{k=1}^{l}(d_{1j}(k)-d_{1\Sigma}(k)+x_{jk})^{2}\\ \vdots\\ {\sum_{k=1}^{l}(d_{ij}(k)-d_{i\Sigma}(k)+x_{jk})^{2}}\\ \vdots\\ {\sum_{k=1}^{l}(d_{nj}(k)-d_{n\Sigma}(k)+x_{jk})^{2}}\\ \end{array}\right)=\left(\begin{array}[]{c}1\\ \vdots\\ 1\\ \vdots\\ 1\\ \end{array}\right),\end{split}$ (17)
    where $d_{ij}(k)$ and $d_{i\Sigma}(k)$ are the $k$th element of the vectors $d_{ij}$ and $d_{i\Sigma}$, respectively. Then, if $l\leq 2n+1$, $S_{2}$ can obtain $g_{j}^{(t)}=\{x_{1j},\cdots,x_{lj}\}$ through solving the joint equation of Eq. (15), Eq. (16) and Eq. (17); Otherwise, $S_{2}$ can launch probabilistic attacks to guess $g_{j}^{(t)}$.
  2. b)
    As a practical attack utilized in [35], $S_{2}$ can register as a benign user $\mathcal{U}_{n+1}$ in ShieldFL scheme and submit its local gradients. This behavior does not violate the honest-but-curious assumption of the server in the threat model.
4.4 Immediate Fix and Its Ineffectiveness
The main reason for the privacy breaches in ShieldFL is that the local gradients $g_{i}^{(t)}=\{x_{i1},\cdots,x_{il}\}$ and the last-iteration aggregated gradient $g_{\Sigma}^{(t-1)}=\{y_{1},\cdots,y_{l}\}$ (or the poisonous gradient baseline $g_{*}^{(t)}=\{z_{1},\cdots,z_{l}\}$) are masked with the same random noise vector, thereby exposing the correlations between these gradients and resulting in the severe consequence that once one of these gradients is leaked, all other gradients can be directly revealed in clear.
Although different random noise vectors can be used on the local gradients, the last-iteration aggregated gradient, and the poisonous gradient baseline to prevent the aforementioned privacy violations, using the same noise vector is inevitable to maintain the correctness of the secure cosine similarity measurement $\mathbf{SecCos}$. In detail, the secure cosine similarity measurement between two gradient ciphertexts $[\![g_{a}]\!]$ and $[\![g_{b}]\!]$ is calculated by first masking the two gradients with random noise vectors $r$ and $\gamma$ to output $[\![\bar{g}_{a}]\!]=[\![g_{a}+r]\!]$ and $[\![\bar{g}_{b}]\!]=[\![g_{b}+\gamma]\!]$, respectively, then decrypting the masked gradients and calculating the cosine similarity $\overline{\cos}_{ab}=\frac{\bar{g}_{a}\odot\bar{g}_{b}}{|\!|\bar{g}_{a}|\!|\cdot|\!|\bar{g}_{b}|\!|}=\frac{(g_{a}+r)\odot(g_{b}+\gamma)}{|\!|g_{a}+r|\!|\cdot|\!|g_{b}+\gamma|\!|}$. However, $\frac{\bar{g}_{a}\odot\bar{g}_{b}}{|\!|\bar{g}_{a}|\!|\cdot|\!|\bar{g}_{b}|\!|}=\frac{g_{a}\odot g_{b}}{|\!|g_{a}|\!|\cdot|\!|g_{b}|\!|}$ holds only when $r=\gamma$, where $\frac{g_{a}\odot g_{b}}{|\!|g_{a}|\!|\cdot|\!|g_{b}|\!|}$ is the correct computation of the cosine similarity between $g_{a}$ and $g_{b}$; Otherwise, if $r\neq\gamma$, the correctness of the calculation will be breached.
To prevent privacy violations and ensure the correctness of computations, we deviate from ShieldFL of employing distinct random noise vectors to obscure gradient ciphertexts. Instead, we introduce a novel method where we mask the computation result of gradient ciphertexts with a random noise vector, as detailed in our new secure computation methods outlined in Section 5.2. This approach leverages a newly constructed two-trapdoor Fully Homomorphic Encryption (FHE) algorithm as the fundamental cryptographic building block.
5 The Enhanced Scheme
In this section, we propose an enhanced PBFL scheme based on the two-server setting and effectively addresses the privacy leakages identified in ShieldFL. Additionally, the enhanced scheme demonstrates improved efficiency in both computational and communication performance.
5.1 Technical Intuition
We empower our enhanced model to resist privacy attacks while ensuring correctness. In ShieldFL’s secure computation method, the server $S_{1}$ employs masking techniques to process the user gradient ciphertexts it receives, then $S_{1}$ and $S_{2}$ interact to compute on the masked ciphertexts, and finally $S_{2}$ partially decrypts the computation results of the masked ciphertexts and sends them back to $S_{1}$. $S_{1}$ obtains the masked computation results and demasks to acquire the final computation plaintext output. While the masking technique is highly efficient, ensuring the correctness of ciphertext computations requires that different ciphertexts be masked with the same random noise, which leads to significant privacy leakage issues.
Our enhanced scheme retains the efficient masking techniques to safeguard the security of interactions between the two servers. However, unlike ShieldFL, each ciphertext in our scheme is masked with randomly selected noise. To ensure correctness of computations on the masked ciphertexts, we categorize operations on ciphertexts into linear (such as addition, subtraction, and scalar multiplication) and non-linear operations (such as multiplication and exponentiation) from the perspective of masking. For linear operations, the computation results of the masked ciphertexts can be correctly demasked; however, non-linear operations introduce errors that are challenging to correct. To address this issue without adding computational burdens, we perform non-linear operations on the original ciphertexts locally at $S_{1}$. After conducting these operations, $S_{1}$ masks the results and sends them to $S_{2}$, which then performs linear operations on the masked ciphertexts and decrypts the results. Consequently, $S_{1}$ can successfully demask $S_{2}$’s decrypted results to obtain the final correct computation output. In our scheme, $S_{1}$ only accesses the final computation results, and $S_{2}$ retains the masked intermediate computation results, thus resisting our privacy attacks.
Building on the technical intuition outlined above, we design two secure computation methods between the servers $S_{1}$ and $S_{2}$: an enhanced secure normalization judgment and an enhanced secure cosine similarity measurement, as introduced in Section 5.2. Notably, our design principle can also be extended to develop other secure computation methods, such as secure Euclidean distance, secure mean, and secure median, by separating the computation into linear and non-linear operations and employing our encryption and masking techniques within the two-server interaction framework.
We improve both computational and communication efficiency in the enhanced scheme. This leap in efficiency stems from a couple of key insights:
(1) Ciphertext computation simplicity: In our enhanced scheme, the secure computation process of non-linear operation on ciphertexts–crucial for operations like computing ${\cos}_{ab}={g}_{a}\odot{g}_{b}$ during $\mathbf{SecCos}([\![g_{a}]\!],[\![g_{b}]\!])$–is streamlined. Unlike previous methods that relied on Paillier’s addition homomorphism and complex masking and unmasking techniques, requiring multiple interactions between the two servers, we construct a two-trapdoor FHE that allows for direct non-linear operations on ciphertexts. This direct approach can significantly reduces both the number of computation steps and communication rounds, making our secure computation methods more efficient than those in ShieldFL. For example, our secure computation methods (outlined in Figs. 3 and 4) only require two rounds of interaction between $S_{1}$ and $S_{2}$, whereas the original scheme requires four rounds, reducing the communication burden by (at least) half.
(2) Cryptographic primitive efficiency: We employ an RLWE-based FHE, specifically the CKKS cryptosystem, to construct our two-trapdoor FHE scheme. This choice is more efficient than the Paillier cryptosystem-based two-trapdoor PHE. The RLWE-based system, with its smaller ciphertext modulus requirements, offers lower computational overhead and communication traffic per step. For instance, at a $128$-bit security level, the two-trapdoor PHE demands a ciphertext modulus $N^{2}$ of at least $3072\times 2$ bits, whereas the RLWE-based FHE can operate with a much smaller $q$ of $56$ bits (with $n_{f}=2048$), significantly reducing the computational cost. We provide a detailed performance comparison in Section 6.2.
In summary, our enhanced scheme introduces significant improvements in both efficiency and security in the realm of secure computation. By employing randomly selected noise for masking and categorizing operations into linear and non-linear types, we effectively address the privacy concerns associated with previous methods. The streamlined process for non-linear operations, facilitated by a two-trapdoor FHE based on the RLWE homomorphic encryption scheme, minimizes the complexity and communication overhead typically required. This innovative approach not only ensures the correctness of the computations but also bolsters resistance to privacy attacks, setting a new benchmark for secure computation methods. Ultimately, our advancements position our scheme as a more efficient and secure alternative to existing solutions like ShieldFL.
5.2 New Secure Computation Methods
Our new secure computation methods consist of an enhanced secure normalization judgment and an enhanced secure cosine similarity measurement. We construct a two-trapdoor FHE scheme to protect the privacy of these two methods. The two-trapdoor FHE is based on RLWE (Ring Learning with Errors) homomorphic encryption (CKKS cryptosystem [37]), mainly containing the following algorithms.
Given a ciphertext $[\![g_{i}]\!]=\{[\![g_{i,1}]\!],\cdots,[\![g_{i,\tau}]\!]\}$ of a local gradient $g_{i}$, where $\tau=\lceil l/n_{f}\rceil$ and $g_{i,j}=\{x_{i,j,1},\cdots,x_{i,j,n_{f}}\}$ for $j\in[1,\tau]$, $S_{1}$ interacts with $S_{2}$ to judge whether the local gradient $g_{i}$ is normalized:
  • $\bullet$
    @$S_{1}$: $S_{1}$ first calculates the component-wise second power of $[\![g_{i}]\!]$ as $[\![g_{i}^{2}]\!]=\{[\![g^{2}_{i,1}]\!],\cdots,[\![g^{2}_{i,\tau}]\!]\}$, where $[\![g^{2}_{i,j}]\!]\leftarrow\mathbf{FHE.Mult}_{evk}([\![g_{i,j}]\!],[\![g_{i,j}]\!])$ for $j\in[1,\tau]$. To protect gradient privacy, $S_{1}$ selects $\tau$ random noise polynomials $r_{i,j}=\{r_{i,j,1},\cdots,r_{i,j,n_{f}}\}\leftarrow\mathcal{R}_{q}$ for ${j\in[1,\tau]}$ and masks $[\![g_{i,j}^{2}]\!]$ with $r_{i,j}$ as $[\![\bar{g}_{i,j}^{2}]\!]=[\![g_{i,j}^{2}+r_{i,j}]\!]\leftarrow\mathbf{FHE.Add}([\![g_{i,j}^{2}]\!],[\![r_{i,j}]\!]).$ Then, $S_{1}$ executes $[\bar{g}_{i,j}^{2}]_{1}\leftarrow\mathbf{FHE.PartDec}_{sk_{1}}([\![\bar{g}_{i,j}^{2}]\!])$ for $j\in[1,\tau]$ and sends $\{[\bar{g}_{i,j}^{2}]_{1},[\![\bar{g}_{i,j}^{2}]\!]\}_{j\in[1,\tau]}$ to $S_{2}$.
  • $\bullet$
    @$S_{2}$: After obtaining the above ciphertexts, $S_{2}$ obtains the masked polynomials $\bar{g}_{i}^{2}=\{\bar{g}_{i,1}^{2},\cdots,\bar{g}_{i,\tau}^{2}\}$ by executing $[\bar{g}_{i,j}^{2}]_{2}\leftarrow\mathbf{FHE.PartDec}_{sk_{2}}([\![\bar{g}_{i,j}^{2}]\!])$ and $\bar{g}_{i,j}^{2}\leftarrow\mathbf{FHE.FullDec}([\![\bar{g}_{i,j}^{2}]\!],[\bar{g}_{i,j}^{2}]_{1},[\bar{g}_{i,j}^{2}]_{2})$ for $j\in[1,\tau]$. According to the multiplicative homomorphic properties of the two-trapdoor FHE, it is held that $\bar{g}_{i,j}^{2}=\{\bar{x}^{2}_{i,j,1},\cdots,\bar{x}^{2}_{i,j,n_{f}}\}$, where $\bar{x}^{2}_{i,j,k}=x^{2}_{i,j,k}+r_{j,k}$ for $k\in[1,n_{f}]$. Then, $S_{2}$ returns $\overline{sum}=\sum_{j=1}^{\tau}\sum_{k=1}^{n_{f}}\bar{x}_{i,j,k}^{2}$ to $S_{1}$.
  • $\bullet$
    @$S_{1}$: On receiving $\overline{sum}$, $S_{1}$ removes the random noises in $\overline{sum}$ by executing $sum=\overline{sum}-\sum_{j=1}^{\tau}\sum_{k=1}^{n_{f}}r_{i,j,k}.$ Then $S_{1}$ determines whether $sum=1$. If yes, it means that $g_{i}$ is normalized; otherwise not.
Fig. 3 Implementation procedure of the enhanced secure normalization judgement method $\mathbf{ESecJudge}([\![g_{i}]\!])$.
  • $\mathbf{FHE.Setup}(1^{\varepsilon^{\prime}})\rightarrow PP$: Given the security parameter $\varepsilon^{\prime}$, the public parameter $PP=\{n_{f},p,q,P,\chi_{s},\chi_{e}\}$ is output, where $n_{f}$ is a ring dimension, $p$ is a plaintext modulus, $q=q(\varepsilon^{\prime})$ is a ciphertext modulus, $P$ is a special modulus, and $\chi_{s},\chi_{e}$ are two probability distributions on the ring $\mathcal{R}$ with small bounds.
  • $\mathbf{FHE.KeyGen}(PP)\rightarrow(sk,pk,evk)$: Given the public parameter $PP$, a secret key $sk\leftarrow\chi_{s}$, a public key $pk\leftarrow(-a\cdot sk+e\bmod q,a)\in\mathcal{R}_{q}^{2}$, and an evaluation key $evk\leftarrow(-a^{\prime}\cdot sk+e^{\prime}+P\cdot sk^{2}\bmod P\cdot q,a^{\prime})\in\mathcal{R}_{P\cdot q}^{2}$ are output, where $a,a^{\prime}\leftarrow\mathcal{R}_{q}:=\mathbb{Z}_{q}/(X^{n_{f}}+1)$ and $e,e^{\prime}\leftarrow\chi_{e}$.
  • $\mathbf{FHE.KeySplit}(sk)\rightarrow(sk_{1},sk_{2})$: Given the secret key $sk$, two secret key shares $sk_{1},sk_{2}$ are generated by randomly dividing $sk$ meeting $sk=sk_{1}+sk_{2}\bmod q$.
  • $\mathbf{FHE.Enc}_{pk}(m)\rightarrow[\![m]\!]$: This algorithm first encodes a float vector $m\in\mathbb{R}^{n_{f}}$ with length $n_{f}$ into an integral coefficient polynomials $m(X)\in\mathcal{R}_{p}$ and then it encrypts $m(X)$ and outputs the ciphertext $[\![m]\!]$. Specifically, the input vector $m$ is first converted to a complex vector $z=(z_{1},\cdots,z_{n_{f}/2})\in\mathbb{C}^{n_{f}/2}$, in which every two elements in $m$ form the real and imaginary parts of a complex number. Then, the complex vector $z$ is encoded as $m(X)=\lfloor\Delta\cdot\phi^{-1}(z)\rceil$, where $\Delta>1$ is a scaling factor and $\phi:R[X]/(X^{n_{f}}+1)\rightarrow\mathbb{C}^{n_{f}/2}$ is a ring isomorphism. Finally, the ciphertext $[\![m]\!]=(m+u\cdot pk_{0}+e_{0}\bmod q,u\cdot pk_{1}+e_{1}\bmod q)\triangleq(c_{0},c_{1})\in\mathcal{R}_{q}^{2}$ is output, where $u\leftarrow\chi_{s}$ and $e_{0},e_{1}\leftarrow\chi_{e}$.
  • $\mathbf{FHE.Add}([\![m_{1}]\!],[\![m_{2}]\!])\rightarrow[\![m]\!]$: Given two ciphertexts $[\![m_{1}]\!],[\![m_{2}]\!]$, the sum $[\![m]\!]=[\![m_{1}]\!]+[\![m_{2}]\!]=[\![m_{1}+m_{2}]\!]$ of the two ciphertexts is output.
  • $\mathbf{FHE.Mult}_{evk}([\![m_{1}]\!],[\![m_{2}]\!])\rightarrow[\![m]\!]$: Given two ciphertexts $[\![m_{1}]\!]=(c_{0},c_{1}),[\![m_{2}]\!]=(c^{\prime}_{0},c^{\prime}_{1})$, let $(d_{0},d_{1},d_{2})=(c_{0}c^{\prime}_{0},c_{0}c^{\prime}_{1}+c^{\prime}_{0}c_{1},c_{1}c^{\prime}_{1})\bmod q$, the product $[\![m]\!]=(d_{0},d_{1})+\lfloor P^{-1}\cdot d_{2}\cdot evk\rceil$ of the two ciphertexts is output.
  • $\mathbf{FHE.PartDec}_{sk_{i}}([\![m]\!])\rightarrow[m]_{i}$: Given a ciphertext $[\![m]\!]=(c_{0},c_{1})$, a partial decryption result $[m]_{i}=c_{1}sk_{i}+e_{i}\bmod q$ is output, where $e_{i}\leftarrow\chi_{e}$.
  • $\mathbf{FHE.FullDec}([\![m]\!],[m]_{1},[m]_{2})\rightarrow m$: Given a ciphertext $[\![m]\!]=(c_{0},c_{1})$ and its partial decryption result pair $([m]_{1},[m]_{2})$, the algorithm first fully decrypts the ciphertext $[\![m]\!]$ as a plaintext polynomial $m(X)=c_{0}+[m]_{1}+[m]_{2}\bmod p\in\mathcal{R}_{p}$, then it decodes $m(X)$ by first computing the complex vector $z=\Delta^{-1}\cdot\phi(m(X))\in\mathbb{C}^{n_{f}/2}$ and then converting $z$ to a float vector $m\in\mathbb{R}^{n_{f}}$. Finally, the algorithm outputs $m$.
Given two normalized ciphertexts $[\![g_{a}]\!]=\{[\![g_{a,1}]\!],\cdots,[\![g_{a,\tau}]\!]\}$ and $[\![g_{b}]\!]=\{[\![g_{b,1}]\!],\cdots,[\![g_{b,\tau}]\!]\}$ of the gradients $g_{a}$ and $g_{b}$, respectively, where $g_{a,j}=\{x_{a,j,1},\cdots,x_{a,j,{n_{f}}}\}$ and $g_{b,j}=\{x_{b,j,1},\cdots,x_{b,j,{n_{f}}}\}$ for $j\in[1,\tau]$, $S_{1}$ interacts with $S_{2}$ to obtain the cosine similarity $\cos_{ab}$ of $g_{a}$ and $g_{b}$:
  • $\bullet$
    @$S_{1}$: On receiving $[\![g_{a}]\!]$ and $[\![g_{b}]\!]$, $S_{1}$ calculates the product of the two ciphertexts as $[\![g_{c,j}]\!]=[\![g_{a,j}]\!]\times[\![g_{b,j}]\!]\leftarrow\mathbf{FHE.Mult}_{evk}([\![g_{a,j}]\!],[\![g_{b,j}]\!])$ for $j\in[1,\tau]$. To protect gradient privacy, $S_{1}$ selects $\tau$ random noise polynomials $r_{j}=\{r_{j,1},\cdots,r_{j,n_{f}}\}\leftarrow\mathcal{R}_{q}$ for ${j\in[1,\tau]}$ and masks $[\![g_{c,j}]\!]$ with $r_{j}$ as $[\![\bar{g}_{c,j}]\!]=[\![g_{c,j}+r_{j}]\!]\leftarrow\mathbf{FHE.Add}([\![g_{c,j}]\!],[\![r_{j}]\!]).$ Then, $S_{1}$ calculates $[\![\bar{g}_{c,\Sigma}]\!]=[\![\sum_{j=1}^{\tau}\bar{g}_{c,j}]\!]\leftarrow\mathbf{FHE.Add}([\![\bar{g}_{c,1}]\!],\cdots,[\![\bar{g}_{c,\tau}]\!])$ and partially decrypts it as $[\bar{g}_{c,\Sigma}]_{1}\leftarrow\mathbf{FHE.PartDec}_{sk_{1}}([\![\bar{g}_{c,\Sigma}]\!])$. $S_{1}$ sends $[\bar{g}_{c,\Sigma}]_{1},[\![\bar{g}_{c,\Sigma}]\!]$ to $S_{2}$.
  • $\bullet$
    @$S_{2}$: After receiving $[\bar{g}_{c,\Sigma}]_{1},[\![\bar{g}_{c,\Sigma}]\!]$, $S_{2}$ obtains $\bar{g}_{c,\Sigma}=\{x_{c,j,1},\cdots,x_{c,j,n_{f}}\}$ by calling $\mathbf{FHE.PartDec}$ and $\mathbf{FHE.FullDec}$. According to the additive and multiplicative homomorphic properties of the two-trapdoor FHE, it is held that $x_{c,j,k}=\sum_{j=1}^{\tau}(x_{a,j,k}x_{b,j,k}+r_{j,k})$. Then, $S_{2}$ calculates the masked cosine similarity $\overline{\cos}_{ab}=\sum_{k=1}^{n_{f}}\sum_{j=1}^{\tau}x_{c,j,k}=\sum_{k=1}^{n_{f}}\sum_{j=1}^{\tau}(x_{a,j,k}x_{b,j,k}+r_{j,k})$ and sends it to $S_{1}$.
  • $\bullet$
    @$S_{1}$: On receiving $\overline{\cos}_{ab}$, $S_{1}$ removes the random noises in $\overline{\cos}_{ab}$ by executing ${\cos}_{ab}=\overline{\cos}_{ab}-\sum_{k=1}^{n_{f}}\sum_{j=1}^{\tau}r_{j,k}$.
Fig. 4 Implementation procedure of the enhanced secure cosine similarity measurement method $\mathbf{ESecCos}([\![g_{a}]\!],[\![g_{b}]\!])$.
Based on the two-trapdoor FHE scheme, we design the enhanced secure normalization judgment method $\mathbf{ESecJudge}$ and the enhanced secure cosine similarity measurement method $\mathbf{ESecCos}$ as Fig. 3 and Fig. 4, respectively.
For $\mathbf{ESecJudge}$, on inputting a gradient ciphertext $[\![g_{i}]\!]=\{[\![g_{i,1}]\!],\cdots,[\![g_{i,\tau}]\!]\}$, $S_{1}$ interacts with $S_{2}$ to judge whether the local gradient $g_{i}$ is normalized, i.e., whether $[\![g_{i,1}^{2}]\!]+\cdots+[\![g_{i,\tau}^{2}]\!]=[\![1]\!]$. In this ciphertext computation, the linear and non-linear operations correspond to additions and multiplications, respectively. We leverage $S_{1}$ to perform multiplication operations locally. The resulting computations are then masked and sent to $S_{2}$ for the addition operations.
Specifically, $S_{1}$ first computes ciphertext multiplications $[\![g^{2}_{i,j}]\!]\leftarrow\mathbf{FHE.Mult}_{evk}([\![g_{i,j}]\!],[\![g_{i,j}]\!])$ for $j\in[1,\tau]$ to obtain $[\![g_{i}^{2}]\!]=\{[\![g^{2}_{i,1}]\!],\cdots,[\![g^{2}_{i,\tau}]\!]\}$. It then masks the result $[\![g_{i}^{2}]\!]$ as $[\![\bar{g}_{i}^{2}]\!]=\{[\![\bar{g}^{2}_{i,1}]\!],\cdots,[\![\bar{g}^{2}_{i,\tau}]\!]\}$, where $[\![\bar{g}_{i,j}^{2}]\!]=[\![g_{i,j}^{2}+r_{i,j}]\!]\leftarrow\mathbf{FHE.Add}([\![g_{i,j}^{2}]\!],[\![r_{i,j}]\!])$ and $r_{i,j}$ is a randomly selected noise polynomial. $S_{1}$ partially decrypts $\{[\![\bar{g}_{i,j}^{2}]\!]\}_{j\in[1,\tau]}$ and sends $\{[\bar{g}_{i,j}^{2}]_{1},[\![\bar{g}_{i,j}^{2}]\!]\}_{j\in[1,\tau]}$ to $S_{2}$. After performing full decryption, $S_{2}$ obtains the masked gradient $\bar{g}_{i}^{2}=\{\bar{g}_{i,1}^{2},\cdots,\bar{g}_{i,\tau}^{2}\}$ and adds all components of the masked gradient (polynomial) to obtain $\overline{sum}=\sum_{j=1}^{\tau}\sum_{k=1}^{n_{f}}\bar{x}_{i,j,k}^{2}$, where $\bar{g}_{i,j}^{2}=\{\bar{x}^{2}_{i,j,1},\cdots,\bar{x}^{2}_{i,j,n_{f}}\}$ and $\bar{x}^{2}_{i,j,k}$ is the $k$th polynomial coefficient of $\bar{g}_{i,j}^{2}$. $S_{2}$ sends $\overline{sum}$ to $S_{1}$, who then obtains $sum=\overline{sum}-\sum_{j=1}^{\tau}\sum_{k=1}^{n_{f}}r_{i,j,k}$ through removing noises. Finally, if $sum=1$, $S_{1}$ determines that $g_{i}$ is normalized and accepts it; otherwise, $S_{1}$ discards the unnormalized $g_{i}$. Fig. 3 illustrates the detailed procedure of $\mathbf{ESecJudge}$.
For $\mathbf{ESecCos}$, on inputting two normalized gradient ciphertexts $[\![g_{a}]\!]=\{[\![g_{a,1}]\!],\cdots,[\![g_{a,\tau}]\!]\}$ and $[\![g_{b}]\!]=\{[\![g_{b,1}]\!],\cdots,[\![g_{b,\tau}]\!]\}$ with $\parallel g_{a}\parallel=1$ and $\parallel g_{b}\parallel=1$, $S_{1}$ interacts with $S_{2}$ to calculate the cosine similarity of $g_{a}$ and $g_{b}$ as $\cos_{ab}=\frac{g_{a}\odot g_{b}}{\parallel g_{a}\parallel\times\parallel g_{b}\parallel}=g_{a,1}\times g_{b,1}+\cdots+g_{a,\tau}\times g_{b,\tau}$. In this ciphertext computation, the linear and non-linear operations correspond to additions and multiplications, respectively. We leverage $S_{1}$ to perform multiplication operations locally. The resulting computations are then masked and sent to $S_{2}$ for the addition operations.
In particular, first, $S_{1}$ performs the ciphertext multiplication and masking to obtain $[\![\bar{g}_{c,\Sigma}]\!]=\sum_{j=1}^{\tau}[\![g_{a,j}\times g_{b,j}+r_{j}]\!]$, where $\{r_{j}\}_{j\in[1,\tau]}$ are randomly selected noise polynomials by $S_{1}$. Then, $S_{2}$ obtains the plaintext $\bar{g}_{c,\Sigma}=\{x_{c,j,1},\cdots,x_{c,j,n_{f}}\}$ by jointly decrypting $[\![\bar{g}_{c,\Sigma}]\!]$ with $S_{1}$ and sends the masked cosine similarity $\overline{\cos}_{ab}=\sum_{k=1}^{n_{f}}\sum_{j=1}^{\tau}x_{c,j,k}$ to $S_{1}$, where $x_{c,j,k}=\sum_{j=1}^{\tau}(g_{a,j,k}\times g_{b,j,k}+r_{j,k})$. Finally, $S_{1}$ obtains $\cos_{ab}$ through removing noises. A more specified $\mathbf{ESecCos}$ procedure is described in Fig. 4.
5.3 The Enhanced Scheme
In our enhanced scheme, a privacy-preserving defense method are devised to fortify against model poisoning attacks, in which the new secure computation methods are utilized to protect the privacy while ensuring the correctness of the computations. Besides, to motivate users to engage in benign training and accurately identify poisonous gradients, inspired by [38], we introduce user credit scores to our Byzantine-tolerance aggregation process of the privacy-preserving defense. Specifically, the enhanced scheme contains four processes detailed below.
  1. 1.
    Setup: In this process, $KC$ generates public parameters/keys for the scheme and $S_{1}$ generates an initial model for the users. Specifically, given a security parameter $\varepsilon^{\prime}$, $KC$ calls $\mathbf{FHE.Setup}$ and $\mathbf{FHE.KeyGen}$ to generate the public parameter $PP$ and a tuple of key $(sk,pk,evk)$. Then, $KC$ calls $\mathbf{FHE.KeySplit}(sk)$ with $n+1$ times to generate $n+1$ pairs of secret key shares $(sk_{1},sk_{2})$ and $\{(sk_{s_{i}},sk_{u_{i}})\}_{i\in[1,n]}$. Finally, $KC$ broadcasts $\{PP,pk,evk\}$, distributes $sk_{1},\{sk_{s_{i}}\}_{i\in[1,n]}$ to $S_{1}$, distributes $sk_{2}$ to $S_{2}$, and distributes $sk_{u_{i}}$ to $U_{i}$. $S_{1}$ selects a random initial model $W^{(0)}$ and sends $W^{(0)}$ to $U_{i}$.
  2. 2.
    Local Training: In each training iteration, each user trains its local model to obtain a local gradient, encrypts the local gradient, and sends the encrypted gradient to $S_{1}$. Specifically, in the $t$th training iteration, each user $U_{i}$ trains its local model $W^{(t)}$ and flats its local gradient as a gradient vector $g_{i}^{(t)}\in\mathbb{R}^{l}$ whose length is $l$. The gradient vector is then normalized and uniformly divided into $\tau=\lceil l/n_{f}\rceil$ vectors $\{g_{i,j}^{(t)}\in\mathbb{R}^{n_{f}}\}_{j\in[1,\tau]}$ with length $n_{f}$. The elements 0s are padded at the end of $g_{i}^{(t)}$ when less than $n_{f}$ elements are left. Then, $U_{i}$ encrypts each vector $g_{i,j}^{(t)}$ as $[\![g_{i,j}^{(t)}]\!]$ by calling $\mathbf{FHE.Enc}$. Finally, the beginning user $U_{i}$ sends its encrypted gradient $[\![g_{i}^{(t)}]\!]=\{[\![g_{i,1}^{(t)}]\!],\cdots,[\![g_{i,\tau}^{(t)}]\!]\}$ to $S_{1}$ and the malicious user $U_{i}\in U^{*}$ may launch poisoning attacks and then send its encrypted poisonous gradients $[\![g_{i*}^{(t)}]\!]=\{[\![g_{i*,1}^{(t)}]\!],\cdots,[\![g_{i*,\tau}^{(t)}]\!]\}$ to $S_{1}$.
  3. 3.
    Privacy-Preserving Defense: This process mainly contains the following three implementations.
    • Normalization judgment. $S_{1}$ determines whether its received local gradient ciphertexts $[\![g_{i}^{(t)}]\!]=$$\{[\![g_{i,1}^{(t)}]\!],\cdots,[\![g_{i,\tau}^{(t)}]\!]\}$ are normalized by calling $\mathbf{ESecJudge}([\![g_{i}^{(t)}]\!])$.
    • Poisonous gradient baseline finding. First, $S_{1}$ calculates the cosine similarity between the last-iteration aggregated gradient ciphertext $[\![g_{\Sigma}^{(t-1)}]\!]$ and each participant’s local gradient ciphertext $[\![g_{i}^{(t)}]\!]$ as
      $\cos_{i\Sigma}\leftarrow\mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!]),\textrm{for\ }i\in[1,n].$
      Then, $S_{1}$ chooses the poisonous baseline $[\![g_{*}^{(t)}]\!]$ from the local gradient ciphertexts $\{[\![g_{i}^{(t)}]\!]\}_{i\in[1,n]}$, whose cosine similarity $\cos_{*\Sigma}$ is the lowest one among $\{\cos_{i\Sigma}\}_{i\in[1,n]}$.
    • Byzantine-tolerance aggregation. To motivate users to engage in benign training and accurately identify poisonous gradients, we construct the Byzantine-tolerance aggregation method as demonstrated below.
      • $S_{1}$ initiates the credit score $cs_{i}=100\%$ for the user $i(i\in[1,N])$.
      • $S_{1}$ computes the cosine similarity between the poisonous baseline $[\![g_{*}^{(t)}]\!]$ and the local gradient ciphertext $[\![g_{i}^{(t)}]\!]$ as $\cos_{i*}\leftarrow\mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!]),\textrm{for\ }i\in[1,n].$ The smaller $\cos_{i*}$ is, the more likely that $[\![g_{i}^{(t)}]\!]$ is a poisonous gradient. Note that $\cos_{i*}\in\mathbb{Z}_{p}$.
      • $S_{1}$ sets the confidence $co_{i}\leftarrow(p-\cos_{i*})\times cs_{i}$ for the gradient $g_{i}^{(t)},i\in(1,N)$ instead of eliminating outliers, to adapt to non-IID datasets among users.
      • $S_{1}$ recomputes all users’ credit scores $cs_{i}=cs_{i}\times\lfloor\frac{co_{i}}{co}\rceil,i\in(1,N)$ for the next iteration, where $co\leftarrow\sum_{j=1}^{n}co_{j}$. The lower the confidence, the more likely the user is malicious, and thus the user’s credit score needs to be reduced accordingly, and on the contrary, it needs to be increased. The credit score contains all history confidences of the user and thus incentives the user to engage in benign training in the whole training process.
      • $S_{1}$ calculates $co_{i}\leftarrow\lfloor\dfrac{co_{i}}{co}\cdot p\rceil$ and weights the local gradients with their respective confidences and then aggregates them to generate a global gradient $[\![g_{\Sigma}^{(t)}]\!]\leftarrow\sum_{i=1}^{n}[\![co_{i}]\!][\![g_{i}^{(t)}]\!]$, where $g_{\Sigma}^{(t)}=\sum_{i=1}^{n}co_{i}g_{i}^{(t)}$.
  4. 4.
    Model Update. $S_{1}$ sends $co,[\![g_{\Sigma}^{(t)}]\!]$ and $[g_{\Sigma}^{(t)}]_{s_{i}}\leftarrow\mathbf{FHE.PartDec}_{sk_{s_{i}}}([\![g_{\Sigma}^{(t)}]\!])$ to $U_{i}$. Then $U_{i}$ calculates the plaintext aggregated gradient $g_{\Sigma}^{(t)}$ by calling $\mathbf{FHE.PartDec}$ and $\mathbf{FHE.FullDec}$ with $sk_{u_{i}}$. Finally, $U_{i}$ updates its local model as $W^{(t+1)}\leftarrow W^{(t)}-co^{-1}\eta^{(t)}g_{\Sigma}^{(t)}$, where $\eta^{(t)}$ is the learning rate.
TABLE III: Computational costs.
Schemes Setup Local Training Privacy-Preserving Defense Model Update Total
$O(\varepsilon+(n+1)\cdot ks$ $O(nlT\cdot train$ $O(4nlT(N\cdot rnd+penc$ $O(3nlT\cdot pdec$ $O(\varepsilon+(n+1)\cdot ks$
ShieldFL $+l\cdot penc$ $+nlT\cdot penc))$ $+9\cdot pdec+padd$ $+nlT\cdot update)$ $+nlT(6N^{4}\cdot mul+40N^{3}\cdot add$
$+(2n+1)l\cdot pdec)$ $+N^{2}\cdot pmul))$ $+8N^{3}\cdot rnd+train+update)$
$O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks$ $O(nlT\cdot train$ $O(n\tau T(3q\cdot rnd+3fenc$ $O(3n\tau T\cdot fdec$ $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks$
Our scheme $+\tau\cdot fenc$ $+n\tau T\cdot fenc))$ $+9fdec+5fadd$ $+nlT\cdot update)$ $+n\tau T(35n_{f}q\cdot mul+59n_{f}q\cdot add$
[…]
TABLE IV: Communication costs.
Schemes Setup Local Training Privacy-Preserving Defense Model Update Total
ShieldFL $O((2n+2)\varepsilon+2nlC_{p})$ $O(nlTC_{p})$ $O(nT(10l+9)C_{p})$ $O(2nlTC_{p})$ $O(2n\varepsilon+13nlTN^{2})$
Our scheme $O((2n+2)n_{f}\varepsilon^{\prime}+2n\tau C_{f})$ $O(n\tau TC_{f})$ $O(nT(2\tau+4)C_{f})$ $O(2n\tau TC_{f})$ $O(2nn_{f}\varepsilon^{\prime}+5n\tau Tq)$
6 Theoretical Analysis
In this section, we delve into an analysis of both the security and complexity of the enhanced scheme.
6.1 Security Analysis
We prove the security of the enhanced scheme based on the security Definition 1.
Definition 1.
Considering a game between an adversary $\mathcal{A}$ and a PPT simulator $\mathcal{S}$, if for any real view REAL of $\mathcal{A}$ and the ideal view IDEAL of $\mathcal{S}$, it holds that
[…]
then the protocol $\prod$ is secure. $\{g_{i}^{(t)}\}_{i\in[1,n]}$ represents all $n$ users’ submitted gradients in the $t$th training iteration.
Theorem 1.
The enhanced scheme can ensure data confidentiality of its privacy-preserving defense process between the two non-colluding and honest-but-curious servers $S_{1}$ and $S_{2}$ over all $n$ users’ encrypted local gradients against a semi-honest adversary $\mathcal{A}$.
Proof.
We use a standard hybrid argument to prove it. For any server $S_{1}$ and $S_{2}$, $\texttt{REAL}_{\mathcal{A}}^{\prod}$ includes all users’ encrypted gradients $\{[\![g_{i}]\!]\}_{i\in[1,n]}$ and intermediate parameters $sum$ and $\cos$ of the Privacy-Preserving Defense protocol $\prod$. A PPT simulator $\mathcal{S}$ analyzes each implementation procedure in the protocol, i.e., $\mathcal{S}$ comprises the implementations of normalization judgment, poisonous gradient baseline finding, and Byzantine-tolerance aggregation. During these procedures, $\mathcal{S}$ makes a series of (polynomially many) subsequent modifications to the random variable REAL, and we prove that any two subsequent random variables are computationally indistinguishable. The detailed proof is given as follows.
$\mathbf{Hyb_{0}}$ This random variable is initialized to be indistinguishable from $\texttt{REAL}_{\mathcal{A}}^{\prod}$ in a real execution process of the protocol.
$\mathbf{Hyb_{1}}$ In this hybrid, the behavior of all simulated users $\{U_{i}\}_{i\in[1,n]}$ are changed that each user $U_{i}$ encrypts its chosen random gradient vector $\tilde{g}_{i}^{(t)}\leftarrow\chi_{g}$ with a public key $pk$ randomly selected by $\mathcal{S}$ to replace the original encrypted gradient vector $[\![g_{i}^{(t)}]\!]$. Since the two-trapdoor FHE is secure under semi-honest adversaries, the two non-colluding honest-but-curious servers cannot distinguish the views between $\{[\![\tilde{g}_{i}^{(t)}]\!]\}_{i\in[1,n]}$ and $\{[\![g_{i}^{(t)}]\!]\}_{i\in[1,n]}$.
$\mathbf{Hyb_{2}}$ In this hybrid, the product of two ciphertexts $[\![g_{i}^{2}]\!]$ in $\mathbf{ESecJudge}$ and $[\![g_{a}]\!]\cdot[\![g_{b}]\!]$ in $\mathbf{ESecCos}$ are substituted with random variables $[\![\chi_{g}^{2}]\!]$ and $[\![\chi_{g}]\!]\cdot[\![\chi_{g}]\!]$, respectively. The security of two-trapdoor FHE and the property of multiplicative homomorphism ensure that this hybrid is indistinguishable from the previous one.
$\mathbf{Hyb_{3}}$ In this hybrid, all masked gradients in the privacy-preserving defense process are substituted. Specifically, we substitute the masked gradients $[\![g_{i,j}^{2}+r_{j}]\!]_{i\in[1,n],j\in[1,\tau]}$ during the executions of $\mathbf{ESecJudge}([\![g_{i}]\!])_{i\in[1,n]}$ in the normalization judgment procedure (resp. the masked gradients $[\![g_{i,j}^{(t)}\cdot g_{\Sigma,j}^{(t-1)}+r_{j}]\!]_{i\in[1,n],j\in[1,\tau]}$ during the executions of $\mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!])_{i\in[1,n]}$ in the poisonous gradient baseline finding procedure and the masked gradients $[\![g_{i,j}^{(t)}\cdot g_{*,j}^{(t)}+r_{j}]\!]_{i\in[1,n],j\in[1,\tau]}$ during the executions of $\mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!])_{i\in[1,n]}$ in the Byzantine-tolerance aggregation procedure) with random variables $[\![\chi_{i,j}]\!]_{i\in[1,n],j\in[1,\tau]}$. The semi-honest security of the two-trapdoor FHE and the non-colluding setting between the two servers guarantee that this hybrid is indistinguishable from the previous one.
$\mathbf{Hyb_{4}}$ In this hybrid, the product of two ciphertexts $[\![g_{i}^{2}]\!]$ in $\mathbf{ESecJudge}$ and $[\![g_{a}]\!]\cdot[\![g_{b}]\!]$ in $\mathbf{ESecCos}$ are separately masked by random polynomial $r_{i}\leftarrow\mathcal{R}_{p}$, which can be perfectly simulated by $\mathcal{S}$. The intermediate parameters, also masked by uniformly random polynomials, maintain a uniform distribution. As addition homomorphism ensures that the distribution of any intermediate variables remains consistent, $\mathcal{A}$ cannot distinguish the output of $\mathcal{S}$ in polynomial time, thus ensuring that this hybrid is identically distributed to the previous one.
$\mathbf{Hyb_{5}}$ In this hybrid, the aggregated gradient $[\![g_{\Sigma}^{(t)}]\!]$ in the Byzantine-tolerance aggregation procedure is calculated using addition homomorphism. The real view of $\mathcal{S}$ is $\texttt{Ideal}_{\mathcal{S}}^{\prod}=\{\{g_{i}{(t)}\}_{i\in[1,n]},$ $sum,\cos,[\![g_{\Sigma}^{(t)}]\!]\}$, where $sum$ and $\cos$ are sets comprising the sums and cosine similarities obtained when executing $\{\mathbf{ESecJudge}([\![g_{i}^{(t)}]\!])\}_{i\in[1,n]}$ and $\{\mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!])\}_{i\in[1,n]}$, respectively. Since the inputs $\{[\![g_{i}^{(t)}]\!]\}_{i\in[1,n]}$ are all encrypted, and the intermediate results are masked by random values, the non-colluding servers involved in calculating $sum$ and $\cos$ cannot compromise participants’ privacy without access to any information about the inputs and intermediate results. Hence, this hybrid is indistinguishable from the previous one.
[…]
6.2 Performance Comparison
Our scheme has lower computational and communication burdens compared with ShieldFL. We compare the complexity of the two schemes in their split processes and the comparison results are shown in Table III and Table IV.
TABLE V: Computational costs of the encryption algorithm in ShieldFL and our scheme.
Schemes Encryption Decryption Homo. Addition Homo. Multiplication
ShieldFL $penc$ $pdec$ $padd$ $pmul$
$O(N(mN\cdot mul+add))$ $O(N((N+1)\cdot mul+add))$ $O(N^{2}\cdot mul)$ $O(2N\cdot rnd+(4mN^{2}+9N^{2}+6N)mul+10N\cdot add)$
Our scheme $fenc$ $fdec$ $fadd$ $fmul$
$O(n_{f}q\cdot(2mul+3add))$ $O(n_{f}q\cdot(mul+3add))$ $O(n_{f}q\cdot add)$ $O(n_{f}q\cdot(5mul+2add))$
TABLE VI: Communication costs of the encryption algorithm in ShieldFL and our scheme
Schemes Homo. multiplication Ciphertext size
ShieldFL $C_{pmul}:O(5N^{2})$ $C_{p}:O(N^{2})$
Our scheme 00 $C_{f}:O(p)$
  • In Setup process of both schemes, KC and $S_{1}$ distribute keys and initial model, respectively. The computation complexities of ShieldFL and our scheme are $O(\varepsilon+(n+1)\cdot ks+l\cdot penc+(2n+1)l\cdot pdec)$ and $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks+\tau\cdot fenc+(2n+1)\tau\cdot fdec)$, respectively, where $ks$, and $penc/fdec$, and $pdec/fdec$ represent the running time when implementing $\mathbf{PHE.KeySplit},$ $\mathbf{PHE.Enc}/$$\mathbf{FHE.Enc},$ and $\mathbf{PHE.Dec}/$$\mathbf{FHE.Dec}$, respectively. The communication complexities for ShieldFL and our scheme are $O((2n+2)\varepsilon+2nlC_{p})$ and $O((2n+2)n_{f}\varepsilon^{\prime}+2n\tau C_{f})$, respectively, where $C_{p}$/$C_{f}$ represents the ciphertext sizes of the two-trapdoor PHE/FHE.
  • In Local Training process of both schemes, $n$ users train their models locally, encrypt their local gradients, and transmit them to the server. The computation complexities of ShieldFL and our scheme are $O(nlT\cdot train+nlT\cdot penc))$ and $O(nlT\cdot train+n\tau T\cdot fenc))$, respectively, where $T$ is the total training iterations. The communication complexities of ShieldFL and our scheme are $O(nlTC_{p})$ and $O(n\tau TC_{f})$, respectively.
  • In Privacy-Preserving Defense process, the computation and communication costs of both schemes mainly come from the protocols $\mathbf{SecJudge}/\mathbf{ESecJudge}$ and $\mathbf{SecCos}/\mathbf{ESecCos}$. For ShieldFL, the computation and communication complexities for executing the two protocols are $O(4nlT(N\cdot rnd+penc+9pdec+padd+N^{2}\cdot pmul))$ and $O(nT(10l+9)C_{p})$, respectively, where $rnd,padd$ and $pmul$ represent the running time of generating a random noise, a homomorphic addition, and a homomorphic multiplication, respectively. For the enhanced scheme, the computation and communication complexities for executing the two protocols are $O(n\tau T(3q\cdot rnd+3fenc+9fdec+5fadd+3fmul))$ and $O(nT(2\tau+4)C_{f})$, respectively, where $fadd/fmul$ represent the running time of a fully homomorphic addition/multiplication.
  • In Model Update process, all users interact with $S_{1}$ to obtain the decrypted aggregated gradient and then update their local models. For ShieldFL, the computation and communication complexities of this process are $O(nlT(3\cdot pdec+update))$ and $O(2nlTC_{p})$, respectively. For our scheme, the computation and communication complexities are $O(3n\tau T\cdot fdec+nlT\cdot update)$ and $O(2n\tau TC_{f})$, respectively.
The total computational costs of ShieldFL and our scheme are $O(\varepsilon+(n+1)\cdot ks+nlT(8penc+36pdec+4padd+4N^{2}\cdot pmul+4N\cdot rnd+train+update)$ and $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks+n\tau T(7fenc+9fdec+5fadd+3fmul+3q\cdot rnd)+nlT\cdot(train+update))$, respectively. Table V presents the computational costs of $penc,pdec,padd,pmul,fenc,fdec,fadd,fmul$ related to general addition $add$ and general multiplication $mul$ operations. Thus, the total computational costs of ShieldFL and our scheme are $O(\varepsilon+(n+1)\cdot ks+nlT(36N^{4}\cdot mul+40N^{3}\cdot add+8N^{3}\cdot rnd+train)$ and $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks+n\tau T(35n_{f}q\cdot mul+59n_{f}q\cdot add+9q\cdot rnd)+nlT\cdot train)$, respectively. The total communication costs of ShieldFL and the enhanced scheme are $O(2n\varepsilon+13nlTC_{p})$ and $O(2nn_{f}\varepsilon^{\prime}+5n\tau TC_{f})$, respectively. Table VI presents the communication costs of homomorphic multiplication operations and the ciphertext sizes of the two schemes. Our scheme’s homomorphic multiplication operation is performed locally, and thus its communication cost is 00. The homomorphic multiplication protocol in ShieldFL requires interactions between the two servers, during which five ciphertexts are transformed. Thus, its communication cost is $O(5N^{2})$. Therefore, the total communication costs of ShieldFL and the enhanced scheme are $O(2n\varepsilon+13nlTN^{2})$ and $O(2nn_{f}\varepsilon^{\prime}+5n\tau Tq)$, respectively.
7 Conclusion
In this paper, we identify security weaknesses in privacy-preserving and Byzantine-robust federated learning (PBFL) schemes with the two-server model and proposes an enhanced PBFL solution. In our enhanced scheme, a Byzantine-tolerant aggregation method is devised to fortify against model poisoning attacks. Additionally, we develop the enhanced secure normalization judgment method and secure cosine similarity measurement method to protect the privacy of the defense process while ensuring the correctness of the computations. Our scheme guarantees privacy preservation and resilience against model poisoning attacks, even in scenarios with heterogeneous datasets that are non-IID (Independently Identically Distributed). Moreover, it demonstrates superior efficiency in computational and communication performance compared to the existing state-of-the-art PBFL schemes.

### v2

@@
 Privacy-Preserving Federated Learning Scheme with Mitigating Model Poisoning Attacks: Vulnerabilities and Countermeasures
-Abstract
-The privacy-preserving federated learning schemes based on the setting of two honest-but-curious and non-colluding servers offer promising solutions in terms of security and efficiency.
-However, our investigation reveals that these schemes still suffer from privacy leakage when considering model poisoning attacks from malicious users.
-Specifically, we demonstrate that the privacy-preserving computation process for defending against model poisoning attacks inadvertently leaks privacy to one of the honest-but-curious servers, enabling it to access users’ gradients in plaintext.
-To address both privacy leakage and model poisoning attacks, we propose an enhanced privacy-preserving and Byzantine-robust federated learning (PBFL) scheme, comprising three components: (1) a two-trapdoor fully homomorphic encryption (FHE) scheme to bolster users’ privacy protection; (2) a novel secure normalization judgment method to preemptively thwart gradient poisoning; and (3) an innovative secure cosine similarity measurement method for detecting model poisoning attacks without compromising data privacy.
-Our scheme guarantees privacy preservation and resilience against model poisoning attacks, even in scenarios with heterogeneous, non-IID (Independently and Identically Distributed) datasets.
-Theoretical analyses substantiate the security and efficiency of our scheme, and extensive experiments corroborate the efficacy of our private attacks.
-Furthermore, the experimental results demonstrate that our scheme accelerates training speed while reducing communication overhead compared to the state-of-the-art PBFL schemes.
+Privacy-Preserving Federated Learning Scheme with Mitigating Model Poisoning Attacks: Vulnerabilities and Countermeasures
+Jiahui Wu, Fucai Luo, Tiecheng Sun, , Haiyan Wang, Weizhe Zhang This work is supported by the Key Program of the Joint Fund of the National Natural Science Foundation of China (Grant No.
+U22A2036) and the Major Key Project of Peng Cheng Laboratory (Grant No.
+PCL2023A06).J.
+Wu, T.
+Sun and H.
+Wang are with the New Network Department, Peng Cheng Laboratory, Shenzhen 518000, China; F.
+Luo is with the School of Computer Science and Technology, Zhejiang Gongshang University, Hangzhou 310018, China; W.
+Zhang is with the School of Cyberspace Science Faculty of Computing, Harbin Institute of Technology, Shenzhen 518055, China and with the New Network Department, Peng Cheng Laboratory, Shenzhen 518000, China.
+E-mail: wujh01@pcl.ac.cn; lfucai@126.com; tiechengsun@126.com; wanghy01@pcl.ac.cn; wzzhang@hit.edu.cn.
+Corresponding author: W.
+Zhang.
 Index Terms:
@@
 In recent years, the US government has been actively promoting research aimed at creating robust and privacy-preserving FL solutions for diverse tasks such as financial crime prevention [7] and pandemic response and forecasting [8].
-Incentives, including substantial prizes totaling \$$\$800,000$, have been offered to spur innovation in this area.
+Incentives, including substantial prizes totaling \$800,000\\$800,000\$ 800 , 000, have been offered to spur innovation in this area.
 Over the years, extensive researches [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have underscored the importance of two critical security properties in FL:
@@
 For instance, researchers propose privacy-preserving and Byzantine-robust FL (PBFL) schemes.
-However, these schemes often suffer from inefficiency [19, 20, 21, 22, 23, 24] or still make privacy leakages [22, 23, 24, 25, 26, 27].
+However, these schemes often suffer from inefficiency [19, 20, 21, 22, 23, 24], still make privacy leakages [22, 23, 24, 25, 26, 27], or not secure for secure multiparty computation (MPC) scenarios like PPFL applications [28, 29].
 Based on server aggregation settings, we classify existing PBFL schemes into three categories: the serverless/decentralized model, the single-server model, and the two-server model.
@@
 Efficiency
-FedSGD [28] $1$
-PrivateFL [9, 10, 11, 12, 13] $0/1/2$
-RobustFL [14, 15, 16, 17, 18] $0/1/2$
-PBFL [22, 23, 24] 00
-PBFL [19, 20, 21] $1$
-PBFL [25, 26, 27] $2$
-Our scheme $2$
-  • Notes. In the column of “Server Setting”, $0,1,$ and $2$ represent the server-less, the single-server, and the two-server models, respectively.
+FedSGD [30] 111
+PrivateFL [9, 10, 11, 12, 13] 0/1/20/1/20 / 1 / 2
+RobustFL [14, 15, 16, 17, 18] 0/1/20/1/20 / 1 / 2
+PBFL [22, 23, 24] 0
+PBFL [19, 20, 21] 111
+PBFL [25, 26, 27, 28, 29] 222
+Our scheme 222
+  • Notes. In the column of “Server Setting”, 0,1,0,1,0 , 1 , and 222 represent the server-less, the single-server, and the two-server models, respectively.
 In the serverless/decentralized model, multiple users collaborate in training a model without any central server.
@@
 These attacks could result in breaches of model privacy through data infringements, hacking, or leaks.
-FL schemes with the two-server model (e.g., [25, 26, 27]) distribute the trust of the central server in the single-server model across two non-colluding servers.
-These schemes present promising solutions by fortifying resilience against single points of failure and ensuring reliable privacy protection among users without necessitating multiple keys/parameters in their privacy-preserving techniques.
-Consequently, they hold the potential to enhance FL performance in terms of computational cost, communication burden, and final model accuracy.
-Nevertheless, current schemes still face challenges in effectively protecting data privacy during their secure defenses against model poisoning attacks.
-In light of these challenges, this paper identifies weaknesses in PBFL schemes with the two-server model and proposes a solution. The main contributions of this paper are threefold:
-(1) We analyze the security of the existing PBFL scheme based on the two-server setting and show that it violates privacy.
+FL schemes with the two-server model (e.g., [25, 26, 27, 28, 29]) distribute the trust assumed in the traditional single-server model across two non-colluding servers.
+These schemes are promising as they eliminate single points of failure and enhance robustness.
+However, even state-of-the-art schemes such as ShieldFL contain a critical privacy vulnerability that can completely undermine their guarantees: the robust aggregation process leaks the relative relationships between clients’ gradients.
+A single compromised client can leverage this leakage to reconstruct other clients’ gradients in clear.
+In ShieldFL [26], each client encrypts its gradient under a unique key, and two servers jointly perform decryption and aggregation.
+To ensure no server directly learns the gradients, ShieldFL masks each gradient with random noise before cosine similarity is computed for robust aggregation.
+However, the same noise is applied to all gradients and to the baseline gradient.
+This design preserves correctness after noise removal, but it also inadvertently reveals the relative offsets between gradients.
+If one client is malicious, it can subtract its own gradient and the shared noise from the noisy similarity results to infer the exact gradients of all other clients, thereby breaching the privacy of the entire system.
+This example highlights a fundamental tension in existing two-server FL schemes: the need to preserve computational correctness (e.g., accurate robust aggregation) often conflicts with data confidentiality.
+Recent works such as [28, 29] employ single-key homomorphic encryption (HE) to protect the confidentiality of data from multiple clients.
+While this approach is effective in some cases, it fails to guarantee confidentiality between clients because all ciphertexts are encrypted under the same key.
+Conversely, schemes based on multi-key HE can protect inter-client confidentiality but face the challenge of performing correct computations across multiple keys.
+For example, existing multi-key HE-based schemes like [25, 26, 27], while offering strong privacy protection, may compromise it in order to ensure computational correctness during model aggregation and poisoning defense.
+In ShieldFL, although encryption and noise masking are employed to hide gradients, the robust aggregation mechanism still leaks information as explained above.
+This vulnerability means that once a client is compromised, the system¡¯s privacy guarantees collapse.
+Therefore, despite advances in secure computation, existing two-server FL schemes still struggle to simultaneously achieve strong data confidentiality and computational correctness in the presence of model poisoning attacks.
+In light of these challenges, this paper identifies weaknesses in PBFL schemes with the two-server model and proposes a solution. The main contributions are threefold:
+(1) We analyze the security of the existing multi-key HE-based PBFL scheme based on the two-server setting and show that it violates privacy.
 Concretely, we present that the main secure computation procedure of the existing two-server-based scheme leaks significant information about the gradients to a semi-honest server.
@@
 Furthermore, it exhibits superior computational and communication efficiency.
-(3) Through experiments, we demonstrate the efficacy of our private attacks on the existing two-server-based scheme, the effectiveness of our defense against model poisoning attacks, and the efficiency of our scheme in terms of computational and communication costs.
+(3) Through experiments, we demonstrate the efficacy of our privacy attacks on the existing two-server-based scheme, the effectiveness of our defense against model poisoning attacks, and the efficiency of our scheme in terms of computational and communication costs.
 The experiment results also show that our scheme outperforms the state-of-the-art PBFL schemes.
@@
 In Section 3, we introduce the system model, the threat model, and the design goals.
-In Section 4, we overview the two-server-based PBFL scheme ShieldFL and present our private attacks to ShieldFL.
+In Section 4, we overview the two-server-based PBFL scheme ShieldFL and present our privacy attacks to ShieldFL.
 To defeat our attacks and achieve design goals, an enhanced scheme is proposed in Section 5.
@@
 $pk,sk,$ $evk$ Public/secret/evaluation key $sk_{i}$ Secret key shares
-[[]],delimited-[]delimited-[][\![$\cdot$]\!],[ [ ⋅ ] ] , $[\cdot]_{i}$ Ciphertext/partially decrypted ciphertext $W^{(t)}$ Model parameter, a flattened vector of all weight matrixes in a learning model
+[[]],[\![$\cdot$]\!],[ [ ⋅ ] ] , $[\cdot]_{i}$ Ciphertext/partially decrypted ciphertext $W^{(t)}$ Model parameter, a flattened vector of all weight matrixes in a learning model
 […]
@@
 Figure 1: The system model
-3.2 Attack Model
-Two types of attackers are considered, each with distinct objectives: Type I attackers aim to learn local gradients to compromise privacy-sensitive information, while Type II attackers seek to poison the model to degrade the final model performance.
-In the system, KC is a fully trusted entity.
-$U_{i\in[1,n]}$ are $n$ curious users, who may be benign or malicious.
-The benign users are Type I attackers who are only curious about other users’ data, while malicious users, in addition to this, may also be manipulated by Type II attackers to launch model poisoning attacks.
-We denote the set of malicious users as $U^{*}$.
-$S_{1}$ and $S_{2}$ are two honest-but-curious, non-colluding servers who honestly perform the given protocol but are curious about users’ privacy.
-The servers are assumed to be vigilant against Type II attackers, as they stand to benefit from an accurate and robust FL model.
-Type I attackers: Both the honest-but-curious servers and the users are regarded as potential Type I attackers.
-A curious server may infer data privacy from local gradients submitted by the users.
-The users may obtain the other uses’ privacy-sensitive information by eavesdropping on the communication channel between the servers and the users.
-Type II attackers: Byzantine adversaries are this type of attacker who manipulate malicious users to deviate from the correct model training by sending poisonous local gradients.
+3.2 Threat Model
+We consider a federated learning system comprising a set of $n$ clients $\{U_{i}\}_{i\in[1,n]}$ and two non-colluding servers $S_{1},S_{2}$.
+The servers collaboratively manage encrypted aggregation and filtering tasks, while the clients periodically upload their local updates.
+We assume the presence of malicious clients and at most one corrupted server, a widely accepted assumption in dual-server frameworks [31, 39].
+To capture different adversarial behaviors, we distinguish between honest-but-curious and malicious corruption types:
+  • Honest-but-curious (semi-honest) parties follow the protocol specification but attempt to infer private information from the data they receive.
+  • Malicious parties may arbitrarily deviate from the protocol, drop or fabricate messages, or tamper with the computation.
+We analyze the following two adversarial scenarios:
+Case 1: Semi-honest server with malicious clients.
+In this setting, both servers behave semi-honest, while some clients may be malicious.
+The semi-honest servers does not deviate from the protocol but attempts to learn sensitive information.
+Under this threat model, our scheme ensures:
+  • The privacy of honest clients’ local updates (e.g., gradients or model parameters) is preserved.
+  • The correctness of server-side computations, including aggregation and filtering, is guaranteed.
+  • Malicious clients cannot compromise the privacy of others or inject invalid influence beyond their own inputs.
+Case 2: Malicious server with malicious clients.
+In this stronger setting, one of the servers may behave maliciously, while the other remains honest-but-curious.
+The corrupted server may alter, drop, or fabricate computation results.
+Nevertheless, as long as the two servers do not collude, our scheme provides:
+  • Privacy protection of honest clients’ updates remains intact due to cryptographic separation and dual control.
+  • However, the correctness of computation cannot be unconditionally guaranteed, as the malicious server may tamper with intermediate results or deviate from the protocol.
+This trade-off is common in dual-server designs [31, 39].
+Similar to SAFEFL [39], our framework is compatible with integrating verifiable computation techniques (e.g., ZKP, MAC-based audits) to ensure correctness even in the presence of a malicious server.
+As verifiability introduces additional complexity and is orthogonal to our core design, we leave its implementation as future work.
+3.3 Design Goals
+The design goals are to ensure security and performance as detailed below.
+  • Confidentiality (Privacy Protection).
+The scheme should be resistant to Type I attackers, ensuring that users’ private datasets cannot be reconstructed by this type of attackers (including other users and servers).
+  • Robustness.
+The scheme should be resistant to Type II attackers.
+It should be capable of detecting model poisoning attacks from malicious users based on users’ encrypted local models and should be able to produce a high-quality global model.
+  • Efficiency. The computational and communication overhead on users and servers should be minimized.
+4 Overview of Prior Work
+In this section, we briefly review the existing two-server-based PBFL scheme, ShieldFL [26], and identify its security vulnerabilities.
+ShieldFL employs a two-trapdoor partial homomorphic encryption (PHE) scheme based on Paillier’s cryptosystem to preserve data privacy.
+The specific construction of the two-trapdoor PHE is provided in Appendix A.
+ShieldFL consists of the following four processes:
+[…]
+2) Local Training: Each benign user $U_{i}$ trains its local model $W^{(t)}$, encrypts its local gradient $g_{i}^{(t)}$, and sends the encrypted gradient $[\![g_{i}^{(t)}]\!]\leftarrow\mathbf{PHE.Enc}_{pk}(g_{i}^{(t)})$ to $S_{1}$.
+A malicious user $U_{i}\in U^{*}$ may launch model poisoning and send encrypted poisonous gradients $[\![g_{i*}^{(t)}]\!]$ to $S_{1}$.
+3) Privacy-Preserving Defense: This process mainly consists of three implementations.
+a) Normalization judgment.
+$S_{1}$ determines whether it received encrypted local gradients $[\![g_{i}^{(t)}]\!]$ are normalized.
+b) Poisonous Gradient Baseline Finding.
+$S_{1}$ finds the gradient $[\![g_{*}^{(t)}]\!]$ (among the users’ local gradients) with the lowest cosine similarity as the baseline of poisonous gradients through implementing a secure cosine similarity measurement $\cos_{i\Sigma}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!])$ in Fig.
+2, where $[\![g_{\Sigma}^{(t-1)}]\!]$ is the aggregated gradient of the last iteration.
+c) Byzantine-Tolerance Aggregation.
+First, $S_{1}$ measures the cosine similarity between the poisonous baseline and each user’s local gradient as $\cos_{i*}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!]).$ Then, $S_{1}$ sets the confidence $co_{i}\leftarrow 1\times deg-\cos_{i*}$ for $g_{i}^{(t)}$ and calculates $co\leftarrow\sum_{i=1}^{n}co_{i},co_{i}=\lfloor\dfrac{co_{i}}{co}\cdot deg\rceil$.
+Finally, $S_{1}$ generates a global gradient ciphertext $[\![g_{\Sigma}^{(t)}]\!]$, where $g_{\Sigma}^{(t)}=\sum_{i=1}^{n}co_{i}g_{i}^{(t)}$.
+4) Model Update.
+$S_{1}$ and $U_{i}$ interact to perform $\mathbf{PHE.PartDec}$ and $\mathbf{PHE.FullDec}$, so as $U_{i}$ obtains the global gradient $g_{\Sigma}^{(t)}$ and updates its local model as $W^{(t+1)}\leftarrow W^{(t)}-\eta^{(t)}g_{\Sigma}^{(t)}$, where $\eta^{(t)}$ is the learning rate of the $t$th training iteration.
 Given two normalized ciphertexts $[\![g_{a}]\!]=\{[\![x_{a_{1}}]\!],\cdots,[\![x_{a_{l}}]\!]\}$ and $[\![g_{b}]\!]=\{[\![x_{b_{1}}]\!],\cdots,[\![x_{b_{l}}]\!]\}$, $S_{1}$ interacts with $S_{2}$ to obtain the cosine similarity $\cos_{ab}$ of $g_{a}$ and $g_{b}$:
@@
 Fig. 2 Implementation procedure of $\mathbf{SecCos}([\![g_{a}]\!],[\![g_{b}]\!])$.
-3.3 Design Goals
-The design goals are to ensure security and performance as detailed below.
-  • Confidentiality (Privacy Protection).
-The scheme should be resistant to Type I attackers, ensuring that users’ private datasets cannot be reconstructed by this type of attackers (including other users and servers).
-  • Robustness.
-The scheme should be resistant to Type II attackers.
-It should be capable of detecting model poisoning attacks from malicious users based on users’ encrypted local models and should be able to produce a high-quality global model.
-  • Efficiency. The computational and communication overhead on users and servers should be minimized.
-4 Security Vulnerabilities of Existing PBFL Schemes
-In this section, we briefly review the existing two-server-based PBFL scheme ShieldFL [26] and identify its security vulnerabilities.
-We begin by discussing the technical background of ShieldFL, followed by outlining its detailed processes.
-Finally, we present the privacy breaches observed in ShieldFL.
-4.1 Technical Background
-ShieldFL utilizes a two-trapdoor partial homomorphic encryption (PHE) scheme based on Paillier’s cryptosystem to protect privacy, which contains five algorithms:
-  • $\mathbf{PHE.KeyGen}(\varepsilon)\rightarrow(pk,sk)$: On input of the security parameter $\varepsilon$, the key generation algorithm first generates two odd primes $p_{1}$ and $p_{2}$ such that $|p_{1}|=|p_{2}|=\varepsilon$ and sets $N=p_{1}\cdot p_{2}$.
-Then, it outputs the public key $pk=(N,1+N)$ and the secret key $sk=\text{lcm}(p_{1}-1,p_{2}-1)$.
-  • $\mathbf{PHE.Enc}_{pk}(m)\rightarrow[\![m]\!]$: On input of a plaintext $m\in\mathbb{Z}_{N}$, the encryption algorithm encrypts $m$ with $pk$ and outputs the ciphertext $[\![m]\!]=(1+N)^{m}\cdot r^{N}\bmod N^{2}$, where $r\in\mathbb{Z}^{*}_{N}$.
-  • $\mathbf{PHE.KeySplit}(sk)\rightarrow(sk_{1},sk_{2})$: Given the secret key $sk$, two secret key shares $sk_{1}$ and $sk_{2}$ are generated through randomly dividing $sk$ meeting $\sum_{i=1}^{2}sk_{i}\equiv 0\bmod sk$ and $\sum_{i=1}^{2}sk_{i}\equiv 1\bmod N.$
-  • $\mathbf{PHE.PartDec}_{sk_{i}}([\![m]\!])\rightarrow[m]_{i}$: Given a ciphertext $[\![m]\!]$, a decryption share $[m]_{i}=[\![m]\!]^{sk_{i}}$ is output through partially decrypting $[\![m]\!]$ with the secret key share $sk_{i}$.
-  • $\mathbf{PHE.FullDec}([m]_{1},[m]_{2})\rightarrow m$: Given the decryption share pair $([m]_{1},[m]_{2})$, the fully decrypted plaintext $m=\frac{(\prod_{i=1}^{2}[m]_{i}\bmod N^{2})-1}{N}\bmod N$ is output.
-More details of the five algorithms can be referred to [26].
-The two-trapdoor PHE scheme meets the additive homomorphism such that $[\![m_{1}]\!]+[\![m_{2}]\!]=[\![m_{1}+m_{2}]\!]$ and $[\![m]\!]^{k}=[\![km]\!]$.
-4.2 ShieldFL Framework
-ShieldFL consists of four processes:
-  1. 1.
-    Setup: In this process, KC and $S_{1}$ distribute keys and an initial model, respectively.
-    • KC generates a public/secret key pair $(pk,sk)$ and calls $\mathbf{PHE.KeySplit}(sk)$ with $n+1$ times to generate $n+1$ pairs of independent secret key shares $(sk_{1},sk_{2})$, $(sk_{s_{1}},sk_{u_{1}}),\cdots,(sk_{s_{n}},sk_{u_{n}})$.
-Then, KC broadcasts $pk$; distributes $sk_{1}$ and $sk_{2}$ to $S_{1}$ and $S_{2}$, respectively; and distributes $\{sk_{s_{i}}\}_{i\in[1,n]}$ and $sk_{u_{i}}$ to $S_{1}$ and $U_{i}$, respectively.
-    • $S_{1}$ releases the encrypted initial model $[\![W^{(0)}]\!]$ and its decryption share $[W^{(0)}]_{s_{i}}\leftarrow\mathbf{PHE.PartDec}_{sk_{s_{i}}}([\![W^{(0)}]\!])$ to $U_{i}$ for $i\in[1,n]$.
-    • Each user $U_{i}$ calculates the decryption share $[W^{(0)}]_{u_{i}}$leftarrow $\mathbf{PHE.PartDec}_{sk_{u_{i}}}([\![W^{(0)}]\!])$ and then obtains the fully decrypted initial model $W^{(0)}$leftarrow $\mathbf{PHE.FullDec}$ $([W^{(0)}]_{s_{i}},[W^{(0)}]_{u_{i}})$ to update its local model.
-  2. 2.
-    Local Training: In each training iteration, each user trains its local model to obtain a local gradient, encrypts the local gradient, and sends the encrypted gradient to $S_{1}$.
-Specifically, in the $t$th iteration, for a begin user $U_{i}$, it trains the local model $W^{(t)}$ and normalizes its local gradient $g_{i}^{(t)}$, then it sends the encrypted normalized gradient $[\![g_{i}^{(t)}]\!]$ to $S_{1}$.
-For a malicious user $U_{i}\in U^{*}$, it launches model poisoning and sends encrypted poisonous gradients $[\![g_{i*}^{(t)}]\!]$ to $S_{1}$.
-  3. 3.
-    Privacy-Preserving Defense: This process mainly consists of the following three implementations.
-    • Normalization judgment.
-$S_{1}$ determines whether it received encrypted local gradients $[\![g_{i}^{(t)}]\!]$ are normalized by implementing a normalization judgment procedure $\mathbf{SecJudge}([\![g_{i}^{(t)}]\!])$.
-    • Poisonous gradient baseline finding.
-$S_{1}$ finds the gradient $[\![g_{*}^{(t)}]\!]$ (among the users’ local gradients) with the lowest cosine similarity as the baseline of poisonous gradients through implementing a secure cosine similarity measurement $\cos_{i\Sigma}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!])$, where $[\![g_{\Sigma}^{(t-1)}]\!]$ is the aggregated gradient of the last iteration.
-    • Byzantine-tolerance aggregation.
-First, $S_{1}$ measures the cosine similarity between the poisonous baseline and each user’s local gradient as $\cos_{i*}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!]).$ Then, $S_{1}$ sets the confidence $co_{i}\leftarrow 1\times deg-\cos_{i*}$ for $g_{i}^{(t)}$ and calculates $co\leftarrow\sum_{i=1}^{n}co_{i},co_{i}=\lfloor\dfrac{co_{i}}{co}\cdot deg\rceil$.
-Finally, $S_{1}$ weights the local gradients with their respective confidences and then aggregates them to generate a global gradient $[\![g_{\Sigma}^{(t)}]\!]$, where $g_{\Sigma}^{(t)}=\sum_{i=1}^{n}co_{i}g_{i}^{(t)}$.
-  4. 4.
-    Model Update.
-$S_{1}$ sends $[\![g_{\Sigma}^{(t)}]\!]$ and $[g_{\Sigma}^{(t)}]_{s_{i}}\leftarrow\mathbf{PHE.PartDec}_{sk_{s_{i}}}([\![g_{\Sigma}^{(t)}]\!])$ to $U_{i}$.
-Then $U_{i}$ decrypts to obtain the aggregated gradient plaintext $g_{\Sigma}^{(t)}$ and updates its local model as $W^{(t+1)}\leftarrow W^{(t)}-\eta^{(t)}g_{\Sigma}^{(t)}$, where $\eta^{(t)}$ is the learning rate of the $t$th training iteration.
-In ShieldFL, privacy leakages primarily occur during the second and third implementations, namely, the poisonous gradient baseline finding and Byzantine-tolerance aggregation, of the privacy-preserving defense process.
-Therefore, we next provide a detailed procedure of the secure computation involved in these two implementations, as illustrated in Fig.
-2: the implementation procedure of the secure cosine similarity measurement method $\mathbf{SecCos}$.
-Further information on secure computations such as $\mathbf{SecJudge}$, please refer to [26].
-For further information on other secure computation procedures such as $\mathbf{SecJudge}$, please refer to [26].
 4.1 Privacy Leakage in ShieldFL
-4.3.1 Poisonous Gradient Baseline Finding
+In ShieldFL, privacy leakages primarily occur during the privacy-preserving defense process.
 In the procedure of poisonous gradient baseline finding, $S_{1}$ interacts with $S_{2}$ to measure the secure cosine similarity $\cos_{i\Sigma}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!])$ for $i\in[1,n]$, so as to find the baseline of poisonous gradient $[\![g_{*}^{(t)}]\!]\in\{[\![g_{i}^{(t)}]\!]\}_{i\in[1,n]}$ with the lowest cosine similarity.
@@
 Then $S_{1}$ partially decrypts $[\![\bar{g}_{i}^{(t)}]\!]$ and $[\![\bar{g}_{\Sigma,i}^{(t)}]\!]$ as $[\bar{g}_{i}^{(t)}]_{1}$ and $[\bar{g}_{\Sigma,i}^{(t)}]_{1}$ and sends $[\![\bar{g}_{i}^{(t)}]\!],[\![\bar{g}_{\Sigma,i}^{(t)}]\!],[\bar{g}_{i}^{(t)}]_{1},[\bar{g}_{\Sigma,i}^{(t)}]_{1}$ to $S_{2}$.
-$\small\left(\begin{array}[]{c}[\![\bar{g}_{1}^{(t)}]\!]\\ \vdots\\ {[\![\bar{g}_{i}^{(t)}]\!]}\\ \vdots\\ {[\![\bar{g}_{n}^{(t)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccc}[\![x_{11}]\!]\cdot[\![r_{11}]\!]&\cdots&[\![x_{1l}]\!]\cdot[\![r_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{i1}]\!]\cdot[\![r_{i1}]\!]}&\cdots&[\![x_{il}]\!]\cdot[\![r_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{n1}]\!]\cdot[\![r_{n1}]\!]}&\cdots&[\![x_{nl}]\!]\cdot[\![r_{nl}]\!]\\ \end{array}\right)$ (1)
-$\small\left(\begin{array}[]{c}[\![\bar{g}_{\Sigma,1}^{(t-1)}]\!]\\ \vdots\\ {[\![\bar{g}_{\Sigma,i}^{(t-1)}]\!]}\\ \vdots\\ {[\![\bar{g}_{\Sigma,n}^{(t-1)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccccc}[\![y_{1}]\!]\cdot[\![r_{11}]\!]&\cdots&[\![y_{l}]\!]\cdot[\![r_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![y_{1}]\!]\cdot[\![r_{i1}]\!]}&\cdots&[\![y_{l}]\!]\cdot[\![r_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![y_{1}]\!]\cdot[\![r_{n1}]\!]}&\cdots&[\![y_{l}]\!]\cdot[\![r_{nl}]\!]\\ \end{array}\right)$ (2)
-Privacy Leakage: Once receiving the ciphertexts $[\![\bar{g}_{i}^{(t)}]\!],[\![\bar{g}_{\Sigma,i}^{(t)}]\!],[\bar{g}_{i}^{(t)}]_{1},[\bar{g}_{\Sigma,i}^{(t)}]_{1}$ from $S_{1}$, the server $S_{2}$ calls $\mathbf{PHE.PartDec}$ and $\mathbf{PHE.FullDec}$ to obtain $\bar{g}_{i}^{(t)}$ and $\bar{g}_{\Sigma,i}^{(t-1)}$ as Eq.
-(3) and Eq.
-(4), respectively.
-$\small\left(\begin{array}[]{c}\bar{g}_{1}^{(t)}\\ \vdots\\ {\bar{g}_{i}^{(t)}}\\ \vdots\\ {\bar{g}_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}x_{11}+r_{11}&\cdots&x_{1l}+r_{1l}\\ \vdots&\cdot&\vdots\\ {x_{i1}+r_{i1}}&\cdots&x_{il}+r_{il}\\ \vdots&\cdot&\vdots\\ {x_{n1}+r_{n1}}&\cdots&x_{nl}+r_{nl}\\ \end{array}\right)$ (3)
-$\small\left(\begin{array}[]{c}\bar{g}_{\Sigma,1}^{(t-1)}\\ \vdots\\ {\bar{g}_{\Sigma,i}^{(t-1)}}\\ \vdots\\ {\bar{g}_{\Sigma,n}^{(t-1)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}y_{1}+r_{11}&\cdots&y_{l}+r_{1l}\\ \vdots&\cdot&\vdots\\ {y_{1}+r_{i1}}&\cdots&y_{l}+r_{il}\\ \vdots&\cdot&\vdots\\ {y_{1}+r_{n1}}&\cdots&y_{l}+r_{nl}\\ \end{array}\right)$ (4)
-Then, $S_{2}$ can learn the difference between two arbitrary local gradients. The difference between the local gradients of the $i$th and $j$th users is calculated as
-$\begin{split}d_{ij}&=g_{i}^{(t)}-g_{j}^{(t)}=\{x_{i1}-x_{j1},\cdots,x_{il}-x_{jl}\}\\ &=\left(\bar{g}_{i}^{(t)}-\bar{g}_{j}^{(t)}\right)-\left(\bar{g}_{\Sigma,i}^{(t-1)}-\bar{g}_{\Sigma,j}^{(t-1)}\right).\end{split}$ (5)
-Besides, $S_{2}$ can learn the difference between the last-iteration aggregated gradient $g_{\Sigma}^{(t-1)}$ and an arbitrary local gradient $g_{i}^{(t)}$:
+$\left(\begin{array}[]{c}[\![\bar{g}_{1}^{(t)}]\!]\\ \vdots\\ {[\![\bar{g}_{i}^{(t)}]\!]}\\ \vdots\\ {[\![\bar{g}_{n}^{(t)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccc}[\![x_{11}]\!]\cdot[\![r_{11}]\!]&\cdots&[\![x_{1l}]\!]\cdot[\![r_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{i1}]\!]\cdot[\![r_{i1}]\!]}&\cdots&[\![x_{il}]\!]\cdot[\![r_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{n1}]\!]\cdot[\![r_{n1}]\!]}&\cdots&[\![x_{nl}]\!]\cdot[\![r_{nl}]\!]\\ \end{array}\right)$ (11)
+$\left(\begin{array}[]{c}[\![\bar{g}_{\Sigma,1}^{(t-1)}]\!]\\ \vdots\\ {[\![\bar{g}_{\Sigma,i}^{(t-1)}]\!]}\\ \vdots\\ {[\![\bar{g}_{\Sigma,n}^{(t-1)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccccc}[\![y_{1}]\!]\cdot[\![r_{11}]\!]&\cdots&[\![y_{l}]\!]\cdot[\![r_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![y_{1}]\!]\cdot[\![r_{i1}]\!]}&\cdots&[\![y_{l}]\!]\cdot[\![r_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![y_{1}]\!]\cdot[\![r_{n1}]\!]}&\cdots&[\![y_{l}]\!]\cdot[\![r_{nl}]\!]\\ \end{array}\right)$ (22)
+Privacy Leakage: $S_{2}$ decrypts the ciphertexts received from $S_{1}$ to obtain $\bar{g}_{i}^{(t)}$ and $\bar{g}_{\Sigma,i}^{(t-1)}$ $(i\in[1,n])$ as follows:
+$\left(\begin{array}[]{c}\bar{g}_{1}^{(t)}\\ \vdots\\ {\bar{g}_{i}^{(t)}}\\ \vdots\\ {\bar{g}_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}x_{11}+r_{11}&\cdots&x_{1l}+r_{1l}\\ \vdots&\cdot&\vdots\\ {x_{i1}+r_{i1}}&\cdots&x_{il}+r_{il}\\ \vdots&\cdot&\vdots\\ {x_{n1}+r_{n1}}&\cdots&x_{nl}+r_{nl}\\ \end{array}\right)$ (33)
+$\left(\begin{array}[]{c}\bar{g}_{\Sigma,1}^{(t-1)}\\ \vdots\\ {\bar{g}_{\Sigma,i}^{(t-1)}}\\ \vdots\\ {\bar{g}_{\Sigma,n}^{(t-1)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}y_{1}+r_{11}&\cdots&y_{l}+r_{1l}\\ \vdots&\cdot&\vdots\\ {y_{1}+r_{i1}}&\cdots&y_{l}+r_{il}\\ \vdots&\cdot&\vdots\\ {y_{1}+r_{n1}}&\cdots&y_{l}+r_{nl}\\ \end{array}\right)$ (44)
+Then, $S_{2}$ can learn the difference between two arbitrary local gradients and the difference between the last-iteration aggregated gradient and an arbitrary local gradient:
+$\begin{split}d_{ij}&=g_{i}^{(t)}-g_{j}^{(t)}=\{x_{i1}-x_{j1},\cdots,x_{il}-x_{jl}\}\\ &=\left(\bar{g}_{i}^{(t)}-\bar{g}_{j}^{(t)}\right)-\left(\bar{g}_{\Sigma,i}^{(t-1)}-\bar{g}_{\Sigma,j}^{(t-1)}\right),i\neq j.\\ \end{split}$ (45)
 $\begin{split}d_{i\Sigma}&=g_{i}^{(t)}-g_{\Sigma}^{(t-1)}=\{x_{i1}-y_{1},\cdots,x_{il}-y_{l}\}\\ &=\bar{g}_{i}^{(t)}-\bar{g}_{\Sigma,i}^{(t-1)}.\end{split}$ (46)
@@
 This compromises privacy significantly [2], as it divulges a considerable amount of information to $S_{2}$, thereby failing to fulfill the confidentiality requirement outlined in [26].
-4.3.2 Byzantine-Tolerance Aggregation
-$S_{1}$ interacts with $S_{2}$ to implement the secure cosine similarity measurement $\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!])$ for $i=1,\cdots,n$, so as to obtain the similarities of all local gradients related to the poisonous gradient baseline $[\![g_{*}^{(t)}]\!]$.
-Specifically, in this procedure, $S_{1}$ selects $l$ random noises $\{\gamma_{i1},\cdots,\gamma_{il}\}$ and masks $[\![g_{i}^{(t)}]\!]=\{[\![x_{i1}]\!],\cdots,[\![x_{il}]\!]\}$ and $[\![g_{*}^{(t)}]\!]=\{[\![z_{1}]\!],\cdots,[\![z_{l}]\!]\}$ as $[\![\tilde{g}_{i}^{(t)}]\!]$ and $[\![\tilde{g}_{*,i}^{(t)}]\!]$ using Eq.
-(8) and Eq.
-(9), respectively.
-Then $S_{1}$ partially decrypts $[\![\tilde{g}_{i}^{(t)}]\!]$ and $[\![\tilde{g}_{*,i}^{(t)}]\!]$ as $[\tilde{g}_{i}^{(t)}]_{1}$ and $[\tilde{g}_{*,i}^{(t)}]_{1}$ and sends $[\![\tilde{g}_{i}^{(t)}]\!],[\![\tilde{g}_{*,i}^{(t)}]\!],[\tilde{g}_{i}^{(t)}]_{1},[\tilde{g}_{*,i}^{(t)}]_{1}$ to $S_{2}$.
-$\small\left(\begin{array}[]{c}[\![\tilde{g}_{1}^{(t)}]\!]\\ \vdots\\ {[\![\tilde{g}_{i}^{(t)}]\!]}\\ \vdots\\ {[\![\tilde{g}_{n}^{(t)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccc}[\![x_{11}]\!]\cdot[\![\gamma_{11}]\!]&\cdots&[\![x_{1l}]\!]\cdot[\![\gamma_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{i1}]\!]\cdot[\![\gamma_{i1}]\!]}&\cdots&[\![x_{il}]\!]\cdot[\![\gamma_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{n1}]\!]\cdot[\![\gamma_{n1}]\!]}&\cdots&[\![x_{nl}]\!]\cdot[\![\gamma_{nl}]\!]\\ \end{array}\right)$ (8)
-$\small\left(\begin{array}[]{c}[\![\tilde{g}_{*,1}^{(t-1)}]\!]\\ \vdots\\ {[\![\tilde{g}_{*,i}^{(t-1)}]\!]}\\ \vdots\\ {[\![\tilde{g}_{*,n}^{(t-1)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccccc}[\![z_{1}]\!]\cdot[\![\gamma_{11}]\!]&\cdots&[\![z_{l}]\!]\cdot[\![\gamma_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![z_{1}]\!]\cdot[\![\gamma_{i1}]\!]}&\cdots&[\![z_{l}]\!]\cdot[\![\gamma_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![z_{1}]\!]\cdot[\![\gamma_{n1}]\!]}&\cdots&[\![z_{l}]\!]\cdot[\![\gamma_{nl}]\!]\\ \end{array}\right)$ (9)
-Privacy Leakage: Upon obtaining the ciphertexts $[\![\tilde{g}_{i}^{(t)}]\!],[\![\tilde{g}_{*,i}^{(t)}]\!],[\tilde{g}_{i}^{(t)}]_{1},[\tilde{g}_{*,i}^{(t)}]_{1}$ from $S_{1}$, the server $S_{2}$ calls $\mathbf{PHE.PartDec}$ and $\mathbf{PHE.FullDec}$ to obtain $\tilde{g}_{i}^{(t)}$ and $\tilde{g}_{*,i}^{(t)}$ as Eq.
-(10) and Eq.
-(11), respectively.
-$\small\left(\begin{array}[]{c}\tilde{g}_{1}^{(t)}\\ \vdots\\ {\tilde{g}_{i}^{(t)}}\\ \vdots\\ {\tilde{g}_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}x_{11}+\gamma_{11}&\cdots&x_{1l}+\gamma_{1l}\\ \vdots&\cdot&\vdots\\ {x_{i1}+\gamma_{i1}}&\cdots&x_{il}+\gamma_{il}\\ \vdots&\cdot&\vdots\\ {x_{n1}+\gamma_{n1}}&\cdots&x_{nl}+\gamma_{nl}\\ \end{array}\right)$ (10)
-$\vspace*{-0.5\baselineskip}\small\left(\begin{array}[]{c}\tilde{g}_{*,1}^{(t)}\\ \vdots\\ {\tilde{g}_{*,i}^{(t)}}\\ \vdots\\ {\tilde{g}_{*,n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}z_{1}+\gamma_{11}&\cdots&z_{l}+\gamma_{1l}\\ \vdots&\cdot&\vdots\\ {z_{1}+\gamma_{i1}}&\cdots&z_{l}+\gamma_{il}\\ \vdots&\cdot&\vdots\\ {z_{1}+\gamma_{n1}}&\cdots&z_{l}+\gamma_{nl}\\ \end{array}\right)$ (11)
-Since $\tilde{g}_{*}^{(t)}$ is the gradient of a certain user, i.e., $\tilde{g}_{*}^{(t)}\in\{\tilde{g}_{1}^{(t)},\cdots,\tilde{g}_{n}^{(t)}\}$, $S_{2}$ can learn which user’s gradient is the poisonous gradient baseline by comparing $\tilde{g}_{*,i}^{(t)}$ and $\tilde{g}_{i}^{(t)}$.
-That is, if $\tilde{g}_{*,k}^{(t)}=\tilde{g}_{k}^{(t)}$, then $S_{2}$ can determine that the gradient of the $k$th user is the poisonous gradient baseline which has the lowest cosine similarity in the $t$th iteration of the model.
-This, to some extent, leaks the $k$th user’s personal information.
-Besides, similar to the analyses in the poisonous gradient baseline finding, in Byzantine-tolerance aggregation, $S_{2}$ can also learn the difference between arbitrary two local gradients from Eq.
-(10) and Eq.
-(11).
-Therefore, $S_{2}$ knows the differences between the gradient of an arbitrary user $i$ and the poisonous gradient baseline, which is
-$\begin{split}d_{i*}&=g_{i}^{(t)}-g_{*}^{(t)}=\{x_{i1}-z_{1},\cdots,x_{il}-z_{l}\}\\ &=\left(\tilde{g}_{i}^{(t)}-\tilde{g}_{l}^{(t)}\right)-\left(\tilde{g}_{*,i}^{(t)}-\tilde{g}_{*,l}^{(t)}\right).\end{split}$ (12)
-Thus, $S_{2}$ knows a “shifted” distribution of all user gradients related to the poisonous gradient baseline through Eq. (13).
-$\small\left(\begin{array}[]{c}g_{1}^{(t)}\\ \vdots\\ {g_{i}^{(t)}}\\ \vdots\\ {g_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{c}d_{1*}+g_{*}^{(t)}\\ \vdots\\ {d_{i*}+g_{*}^{(t)}}\\ \vdots\\ {d_{n*}+g_{*}^{(t)}}\\ \end{array}\right).$ (13)
-In addition, $S_{2}$ can learn the difference between the poisonous gradient baseline and the last-iteration aggregated gradient:
-$\begin{split}d_{*\Sigma}&=g_{*}^{(t)}-g_{\Sigma}^{(t-1)}=\{z_{1}-y_{1},\cdots,z_{l}-y_{l}\}\\ &=\left(\tilde{g}_{*,i}^{(t)}-\bar{g}_{\Sigma,i}^{(t-1)}\right)-\left(\tilde{g}_{i}^{(t)}-\bar{g}_{i}^{(t)}\right).\end{split}$ (14)
-Thus, the correlation between the poisonous gradient baseline $g_{*}^{(t)}$ and the last-iteration aggregated gradient $g_{\Sigma}^{(t-1)}$ is leaked to $S_{2}$, which is $g_{*}^{(t)}=g_{\Sigma}^{(t-1)}+d_{*\Sigma}$ according to Eq.
-(14).
-4.3.3 Gradient Leakage in Clear
-Based on the privacy breaches mentioned above, the local gradients can be directly leaked in plaintext, thus breaking the privacy of the whole ShieldFL scheme.
-Specifically, once $S_{2}$ knows one of the gradients in $\{g_{*}^{(t)},g_{\Sigma}^{(t-1)},\{g_{i}^{(t)}\}_{i\in[1,n]}\}$, it can directly obtain all other users’ local gradients in clear using Eq.
-(7) or Eq.
-(13).
-We present that $S_{2}$ can obtain one of these gradients using the following methods.
-  1. a)
-    Since $\{g_{*}^{(t)},g_{\Sigma}^{(t-1)},\{g_{i}^{(t)}\}_{i\in[1,n]}\}$ are all normalized gradients, and the difference between any two of these gradients is a known value, it is sufficient to $S_{2}$ to guess a gradient in $\{g_{*}^{(t)},g_{\Sigma}^{(t-1)},\{g_{i}^{(t)}\}_{i\in[1,n]}\}$.
-For example, $S_{2}$ knows that
-    $|\!|g_{j}^{(t)}|\!|={{x}_{j1}^{2}}+\cdots+{x}_{jl}^{2}=1.$ (15)
-    According to Eq. (7), $S_{2}$ knows that
-    $\tiny\left(\begin{array}[]{c}|\!|g_{1}^{(t)}|\!|\\ \vdots\\ {|\!|g_{i}^{(t)}|\!|}\\ \vdots\\ {|\!|g_{n}^{(t)}|\!|}\\ \end{array}\right)=\left(\begin{array}[]{c}|\!|d_{1j}+g_{j}^{(t)}|\!|\\ \vdots\\ {|\!|d_{ij}+g_{j}^{(t)}|\!|}\\ \vdots\\ {|\!|d_{nj}+g_{j}^{(t)}|\!|}\\ \end{array}\right)=\left(\begin{array}[]{c}\sum_{k=1}^{l}(d_{1j}(k)+x_{jk})^{2}\\ \vdots\\ {\sum_{k=1}^{l}(d_{ij}(k)+x_{jk})^{2}}\\ \vdots\\ {\sum_{k=1}^{l}(d_{nj}(k)+x_{jk})^{2}}\\ \end{array}\right)=\left(\begin{array}[]{c}1\\ \vdots\\ 1\\ \vdots\\ 1\\ \end{array}\right),$ (16)
-    $\tiny\begin{split}&\left(\begin{array}[]{c}|\!|g_{\Sigma}^{(t-1)}|\!|\\ \vdots\\ {|\!|g_{\Sigma}^{(t-1)}|\!|}\\ \vdots\\ {|\!|g_{\Sigma}^{(t-1)}|\!|}\\ \end{array}\right)=\left(\begin{array}[]{c}|\!|d_{1j}+g_{j}^{(t)}-d_{1\Sigma}|\!|\\ \vdots\\ {|\!|d_{ij}+g_{j}^{(t)}-d_{i\Sigma}|\!|}\\ \vdots\\ {|\!|d_{nj}+g_{j}^{(t)}-d_{n\Sigma}|\!|}\\ \end{array}\right)\\ =&\left(\begin{array}[]{c}\sum_{k=1}^{l}(d_{1j}(k)-d_{1\Sigma}(k)+x_{jk})^{2}\\ \vdots\\ {\sum_{k=1}^{l}(d_{ij}(k)-d_{i\Sigma}(k)+x_{jk})^{2}}\\ \vdots\\ {\sum_{k=1}^{l}(d_{nj}(k)-d_{n\Sigma}(k)+x_{jk})^{2}}\\ \end{array}\right)=\left(\begin{array}[]{c}1\\ \vdots\\ 1\\ \vdots\\ 1\\ \end{array}\right),\end{split}$ (17)
-    where $d_{ij}(k)$ and $d_{i\Sigma}(k)$ are the $k$th element of the vectors $d_{ij}$ and $d_{i\Sigma}$, respectively.
-Then, if $l\leq 2n+1$, $S_{2}$ can obtain $g_{j}^{(t)}=\{x_{1j},\cdots,x_{lj}\}$ through solving the joint equation of Eq.
-(15), Eq.
-(16) and Eq.
-(17); Otherwise, $S_{2}$ can launch probabilistic attacks to guess $g_{j}^{(t)}$.
-  2. b)
-    As a practical attack utilized in [35], $S_{2}$ can register as a benign user $\mathcal{U}_{n+1}$ in ShieldFL scheme and submit its local gradients.
-This behavior does not violate the honest-but-curious assumption of the server in the threat model.
+The privacy leakage issue in the Byzantine-tolerant aggregation phase is similar to that in the poisonous gradient baseline finding process.
+For brevity, we present its details in Appendix B.
+In addition, Appendix B further provides a gradient reconstruction analysis, showing how the leaked information from both phases can be used to recover the complete plaintext gradients.
 4.2 Immediate Fix and Its Ineffectiveness
-The main reason for the privacy breaches in ShieldFL is that the local gradients $g_{i}^{(t)}=\{x_{i1},\cdots,x_{il}\}$ and the last-iteration aggregated gradient $g_{\Sigma}^{(t-1)}=\{y_{1},\cdots,y_{l}\}$ (or the poisonous gradient baseline $g_{*}^{(t)}=\{z_{1},\cdots,z_{l}\}$) are masked with the same random noise vector, thereby exposing the correlations between these gradients and resulting in the severe consequence that once one of these gradients is leaked, all other gradients can be directly revealed in clear.
-Although different random noise vectors can be used on the local gradients, the last-iteration aggregated gradient, and the poisonous gradient baseline to prevent the aforementioned privacy violations, using the same noise vector is inevitable to maintain the correctness of the secure cosine similarity measurement $\mathbf{SecCos}$.
+The main reason for the privacy breaches in ShieldFL is that the local gradients $g_{i}^{(t)}$ and the last-iteration aggregated gradient $g_{\Sigma}^{(t-1)}$ (or the poisonous gradient baseline $g_{*}^{(t)}$) are masked with the same random noise vector, thereby exposing the correlations between these gradients and resulting in the severe consequence that once one of these gradients is leaked, all other gradients can be directly revealed in clear.
+Although different random noise vectors can be used to prevent the aforementioned privacy violations, using the same noise vector is inevitable to maintain the correctness of $\mathbf{SecCos}$.
 In detail, the secure cosine similarity measurement between two gradient ciphertexts $[\![g_{a}]\!]$ and $[\![g_{b}]\!]$ is calculated by first masking the two gradients with random noise vectors $r$ and $\gamma$ to output $[\![\bar{g}_{a}]\!]=[\![g_{a}+r]\!]$ and $[\![\bar{g}_{b}]\!]=[\![g_{b}+\gamma]\!]$, respectively, then decrypting the masked gradients and calculating the cosine similarity $\overline{\cos}_{ab}=\frac{\bar{g}_{a}\odot\bar{g}_{b}}{|\!|\bar{g}_{a}|\!|\cdot|\!|\bar{g}_{b}|\!|}=\frac{(g_{a}+r)\odot(g_{b}+\gamma)}{|\!|g_{a}+r|\!|\cdot|\!|g_{b}+\gamma|\!|}$.
-However, $\frac{\bar{g}_{a}\odot\bar{g}_{b}}{|\!|\bar{g}_{a}|\!|\cdot|\!|\bar{g}_{b}|\!|}=\frac{g_{a}\odot g_{b}}{|\!|g_{a}|\!|\cdot|\!|g_{b}|\!|}$ holds only when $r=\gamma$, where $\frac{g_{a}\odot g_{b}}{|\!|g_{a}|\!|\cdot|\!|g_{b}|\!|}$ is the correct computation of the cosine similarity between $g_{a}$ and $g_{b}$; Otherwise, if $r\neq\gamma$, the correctness of the calculation will be breached.
+However, $\frac{\bar{g}_{a}\odot\bar{g}_{b}}{|\!|\bar{g}_{a}|\!|\cdot|\!|\bar{g}_{b}|\!|}=\frac{g_{a}\odot g_{b}}{|\!|g_{a}|\!|\cdot|\!|g_{b}|\!|}$ holds only when $r=\gamma$, where $\frac{g_{a}\odot g_{b}}{|\!|g_{a}|\!|\cdot|\!|g_{b}|\!|}$ is the correct computation of the cosine similarity between $g_{a}$ and $g_{b}$; Otherwise, if $r\neq\gamma$, the correctness will be breached.
 To prevent privacy violations and ensure the correctness of computations, we deviate from ShieldFL of employing distinct random noise vectors to obscure gradient ciphertexts.
@@
 5.1 Technical Intuition
-We empower our enhanced model to resist privacy attacks while ensuring correctness.
-In ShieldFL’s secure computation method, the server $S_{1}$ employs masking techniques to process the user gradient ciphertexts it receives, then $S_{1}$ and $S_{2}$ interact to compute on the masked ciphertexts, and finally $S_{2}$ partially decrypts the computation results of the masked ciphertexts and sends them back to $S_{1}$.
+We empower our enhanced scheme to resist privacy attacks while ensuring computation correctness.
+In ShieldFL’s secure computation methods, the server $S_{1}$ applies a masking technique to the received user gradient ciphertexts.
+Then $S_{1}$ and $S_{2}$ interact to compute on the masked ciphertexts, and finally $S_{2}$ partially decrypts the computation results of the masked ciphertexts and sends them back to $S_{1}$.
 $S_{1}$ obtains the masked computation results and demasks to acquire the final computation plaintext output.
-While the masking technique is highly efficient, ensuring the correctness of ciphertext computations requires that different ciphertexts be masked with the same random noise, which leads to significant privacy leakage issues.
-Our enhanced scheme retains the efficient masking techniques to safeguard the security of interactions between the two servers.
-However, unlike ShieldFL, each ciphertext in our scheme is masked with randomly selected noise.
-To ensure correctness of computations on the masked ciphertexts, we categorize operations on ciphertexts into linear (such as addition, subtraction, and scalar multiplication) and non-linear operations (such as multiplication and exponentiation) from the perspective of masking.
-For linear operations, the computation results of the masked ciphertexts can be correctly demasked; however, non-linear operations introduce errors that are challenging to correct.
-To address this issue without adding computational burdens, we perform non-linear operations on the original ciphertexts locally at $S_{1}$.
-After conducting these operations, $S_{1}$ masks the results and sends them to $S_{2}$, which then performs linear operations on the masked ciphertexts and decrypts the results.
-Consequently, $S_{1}$ can successfully demask $S_{2}$’s decrypted results to obtain the final correct computation output.
-In our scheme, $S_{1}$ only accesses the final computation results, and $S_{2}$ retains the masked intermediate computation results, thus resisting our privacy attacks.
+Although the masking technique is highly efficient, ensuring the correctness of computation requires that all ciphertexts be masked using the same random noise.
+This introduces strong correlations among masked values: if any plaintext is leaked, all correlated gradients can be inferred, thereby violating privacy.
+A straightforward solution, using independent noise for each ciphertext, would eliminate these correlations but compromise computation correctness.
+To simultaneously ensure both security and correctness, we fundamentally redesign the secure computation paradigm in our enhanced scheme.
+Specifically, we introduce a novel masking protocol build upon a newly constructed two-trapdoor FHE scheme.
+Instead of masking the inputs, we mask the computation results using different random vectors.
+This eliminates the need for shared noise and preserves correctness.
+To handle ciphertext operations under masking, we classify them into linear (e.g., addition, subtraction, scalar multiplication) and non-linear (e.g., multiplication, exponentiation) categories.
+Linear operations on masked ciphertexts yield correctly demaskable results.
+In contrast, non-linear operations introduce complex noise that is difficult to eliminate.
+To mitigate this without adding computational overhead, we perform non-linear operations directly on the original ciphertexts at $S_{1}$.
+The results are then masked and sent to $S_{2}$, which performs the remaining linear operations and partial decryption.
+Finally, $S_{1}$ removes the mask to obtain the correct plaintext output.
+In this design, $S_{1}$ only observes the final computation results, while $S_{2}$ only accesses masked intermediate values.
+This separation effectively defends against the proposed privacy attacks.
 Building on the technical intuition outlined above, we design two secure computation methods between the servers $S_{1}$ and $S_{2}$: an enhanced secure normalization judgment and an enhanced secure cosine similarity measurement, as introduced in Section 5.2.
@@
 We improve both computational and communication efficiency in the enhanced scheme. This leap in efficiency stems from a couple of key insights:
-(1) Ciphertext computation simplicity: In our enhanced scheme, the secure computation process of non-linear operation on ciphertexts–crucial for operations like computing ${\cos}_{ab}={g}_{a}\odot{g}_{b}$ during $\mathbf{SecCos}([\![g_{a}]\!],[\![g_{b}]\!])$–is streamlined.
-Unlike previous methods that relied on Paillier’s addition homomorphism and complex masking and unmasking techniques, requiring multiple interactions between the two servers, we construct a two-trapdoor FHE that allows for direct non-linear operations on ciphertexts.
-This direct approach can significantly reduces both the number of computation steps and communication rounds, making our secure computation methods more efficient than those in ShieldFL.
+(1) Ciphertext computation simplicity: In our enhanced scheme, the secure computation process of non-linear operations on ciphertexts – crucial for operations like computing ${\cos}_{ab}=g_{a}\odot g_{b}$ during $\mathbf{SecCos}([\![g_{a}]\!],[\![g_{b}]\!])$ – is streamlined.
+Unlike previous methods that relied on Paillier’s addition homomorphism and complex masking/unmasking techniques, requiring multiple interactions between the two servers, we construct a two-trapdoor FHE that allows for direct non-linear operations on ciphertexts.
+This direct approach significantly reduces both the number of computation steps and the communication rounds, making our secure computation methods more efficient than those in ShieldFL.
 For example, our secure computation methods (outlined in Figs.
@@
 The RLWE-based system, with its smaller ciphertext modulus requirements, offers lower computational overhead and communication traffic per step.
-For instance, at a $128$-bit security level, the two-trapdoor PHE demands a ciphertext modulus $N^{2}$ of at least $3072\times 2$ bits, whereas the RLWE-based FHE can operate with a much smaller $q$ of $56$ bits (with $n_{f}=2048$), significantly reducing the computational cost.
+For instance, at a 128128128-bit security level, the two-trapdoor PHE demands a ciphertext modulus $N^{2}$ of at least 3072×23072$\times$ 23072 × 2 bits, whereas the RLWE-based FHE can operate with a much smaller $q$ of 565656 bits (with $n_{f}=2048$), significantly reducing the computational cost.
 We provide a detailed performance comparison in Section 6.3.
-In summary, our enhanced scheme introduces significant improvements in both efficiency and security in the realm of secure computation.
+(3) No additional trade-offs: Importantly, these improvements do not introduce extra limitations.
+Our enhanced scheme retains the same flexibility and scalability as ShieldFL because both adopt a two-trapdoor architecture and the same key distribution mechanism.
+Although CKKS introduces negligible approximation error due to its encoding and encryption, such error is naturally tolerated by stochastic machine learning models and does not impact convergence or accuracy, as verified by our experimental results in Section LABEL:sec:experiment.
+In summary, our enhanced scheme introduces significant improvements in both efficiency and security without sacrificing flexibility, accuracy, or scalability.
 By employing randomly selected noise for masking and categorizing operations into linear and non-linear types, we effectively address the privacy concerns associated with previous methods.
@@
   • $\mathbf{FHE.KeySplit}(sk)\rightarrow(sk_{1},sk_{2})$: Given the secret key $sk$, two secret key shares $sk_{1},sk_{2}$ are generated by randomly dividing $sk$ meeting $sk=sk_{1}+sk_{2}\bmod q$.
-  • $\mathbf{FHE.Enc}_{pk}(m)\rightarrow[\![m]\!]$: This algorithm first encodes a float vector $m\in\mathbb{R}^{n_{f}}$ with length $n_{f}$ into an integral coefficient polynomials $m(X)\in\mathcal{R}_{p}$ and then it encrypts $m(X)$ and outputs the ciphertext $[\![m]\!]$.
-Specifically, the input vector $m$ is first converted to a complex vector $z=(z_{1},\cdots,z_{n_{f}/2})\in\mathbb{C}^{n_{f}/2}$, in which every two elements in $m$ form the real and imaginary parts of a complex number.
-Then, the complex vector $z$ is encoded as $m(X)=\lfloor\Delta\cdot\phi^{-1}(z)\rceil$, where $\Delta>1$ is a scaling factor and $\phi:R[X]/(X^{n_{f}}+1)\rightarrow\mathbb{C}^{n_{f}/2}$ is a ring isomorphism.
-Finally, the ciphertext $[\![m]\!]=(m+u\cdot pk_{0}+e_{0}\bmod q,u\cdot pk_{1}+e_{1}\bmod q)\triangleq(c_{0},c_{1})\in\mathcal{R}_{q}^{2}$ is output, where $u\leftarrow\chi_{s}$ and $e_{0},e_{1}\leftarrow\chi_{e}$.
+[…]
   • $\mathbf{FHE.Add}([\![m_{1}]\!],[\![m_{2}]\!])\rightarrow[\![m]\!]$: Given two ciphertexts $[\![m_{1}]\!],[\![m_{2}]\!]$, the sum $[\![m]\!]=[\![m_{1}]\!]+[\![m_{2}]\!]=[\![m_{1}+m_{2}]\!]$ of the two ciphertexts is output.
@@
   • $\mathbf{FHE.PartDec}_{sk_{i}}([\![m]\!])\rightarrow[m]_{i}$: Given a ciphertext $[\![m]\!]=(c_{0},c_{1})$, a partial decryption result $[m]_{i}=c_{1}sk_{i}+e_{i}\bmod q$ is output, where $e_{i}\leftarrow\chi_{e}$.
-  • $\mathbf{FHE.FullDec}([\![m]\!],[m]_{1},[m]_{2})\rightarrow m$: Given a ciphertext $[\![m]\!]=(c_{0},c_{1})$ and its partial decryption result pair $([m]_{1},[m]_{2})$, the algorithm first fully decrypts the ciphertext $[\![m]\!]$ as a plaintext polynomial $m(X)=c_{0}+[m]_{1}+[m]_{2}\bmod p\in\mathcal{R}_{p}$, then it decodes $m(X)$ by first computing the complex vector $z=\Delta^{-1}\cdot\phi(m(X))\in\mathbb{C}^{n_{f}/2}$ and then converting $z$ to a float vector $m\in\mathbb{R}^{n_{f}}$.
-Finally, the algorithm outputs $m$.
+[…]
 Given two normalized ciphertexts $[\![g_{a}]\!]=\{[\![g_{a,1}]\!],\cdots,[\![g_{a,\tau}]\!]\}$ and $[\![g_{b}]\!]=\{[\![g_{b,1}]\!],\cdots,[\![g_{b,\tau}]\!]\}$ of the gradients $g_{a}$ and $g_{b}$, respectively, where $g_{a,j}=\{x_{a,j,1},\cdots,x_{a,j,{n_{f}}}\}$ and $g_{b,j}=\{x_{b,j,1},\cdots,x_{b,j,{n_{f}}}\}$ for $j\in[1,\tau]$, $S_{1}$ interacts with $S_{2}$ to obtain the cosine similarity $\cos_{ab}$ of $g_{a}$ and $g_{b}$:
@@
 4.
-5.3 The Enhanced Scheme
+5.3 Construction of Our PBFL Scheme
 In our enhanced scheme, a privacy-preserving defense method are devised to fortify against model poisoning attacks, in which the new secure computation methods are utilized to protect the privacy while ensuring the correctness of the computations.
@@
 Specifically, the enhanced scheme contains four processes detailed below.
-  1. 1.
-    Setup: In this process, $KC$ generates public parameters/keys for the scheme and $S_{1}$ generates an initial model for the users.
-Specifically, given a security parameter $\varepsilon^{\prime}$, $KC$ calls $\mathbf{FHE.Setup}$ and $\mathbf{FHE.KeyGen}$ to generate the public parameter $PP$ and a tuple of key $(sk,pk,evk)$.
-Then, $KC$ calls $\mathbf{FHE.KeySplit}(sk)$ with $n+1$ times to generate $n+1$ pairs of secret key shares $(sk_{1},sk_{2})$ and $\{(sk_{s_{i}},sk_{u_{i}})\}_{i\in[1,n]}$.
-Finally, $KC$ broadcasts $\{PP,pk,evk\}$, distributes $sk_{1},\{sk_{s_{i}}\}_{i\in[1,n]}$ to $S_{1}$, distributes $sk_{2}$ to $S_{2}$, and distributes $sk_{u_{i}}$ to $U_{i}$.
-$S_{1}$ selects a random initial model $W^{(0)}$ and sends $W^{(0)}$ to $U_{i}$.
-  2. 2.
-    Local Training: In each training iteration, each user trains its local model to obtain a local gradient, encrypts the local gradient, and sends the encrypted gradient to $S_{1}$.
+[…]
+2) Local Training: In each training iteration, each user trains its local model to obtain a local gradient, encrypts the local gradient, and sends the encrypted gradient to $S_{1}$.
 Specifically, in the $t$th training iteration, each user $U_{i}$ trains its local model $W^{(t)}$ and flats its local gradient as a gradient vector $g_{i}^{(t)}\in\mathbb{R}^{l}$ whose length is $l$.
@@
 Finally, the beginning user $U_{i}$ sends its encrypted gradient $[\![g_{i}^{(t)}]\!]=\{[\![g_{i,1}^{(t)}]\!],\cdots,[\![g_{i,\tau}^{(t)}]\!]\}$ to $S_{1}$ and the malicious user $U_{i}\in U^{*}$ may launch poisoning attacks and then send its encrypted poisonous gradients $[\![g_{i*}^{(t)}]\!]=\{[\![g_{i*,1}^{(t)}]\!],\cdots,[\![g_{i*,\tau}^{(t)}]\!]\}$ to $S_{1}$.
-  3. 3.
-    Privacy-Preserving Defense: This process mainly contains the following three implementations.
+3) Privacy-Preserving Defense: This process mainly contains the following three implementations.
   • Normalization Judgment.
-$S_{1}$ determines whether its received local gradient ciphertexts $[\![g_{i}^{(t)}]\!]=$$\{[\![g_{i,1}^{(t)}]\!],\cdots,[\![g_{i,\tau}^{(t)}]\!]\}$ are normalized by calling $\mathbf{ESecJudge}([\![g_{i}^{(t)}]\!])$.
-    • Poisonous gradient baseline finding.
-First, $S_{1}$ calculates the cosine similarity between the last-iteration aggregated gradient ciphertext $[\![g_{\Sigma}^{(t-1)}]\!]$ and each participant’s local gradient ciphertext $[\![g_{i}^{(t)}]\!]$ as
-      $\cos_{i\Sigma}\leftarrow\mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!]),\textrm{for\ }i\in[1,n].$
-      Then, $S_{1}$ chooses the poisonous baseline $[\![g_{*}^{(t)}]\!]$ from the local gradient ciphertexts $\{[\![g_{i}^{(t)}]\!]\}_{i\in[1,n]}$, whose cosine similarity $\cos_{*\Sigma}$ is the lowest one among $\{\cos_{i\Sigma}\}_{i\in[1,n]}$.
-    • Byzantine-tolerance aggregation.
-To motivate users to engage in benign training and accurately identify poisonous gradients, we construct the Byzantine-tolerance aggregation method as demonstrated below.
-      • $S_{1}$ initiates the credit score $cs_{i}=100\%$ for the user $i(i\in[1,N])$.
-      • $S_{1}$ computes the cosine similarity between the poisonous baseline $[\![g_{*}^{(t)}]\!]$ and the local gradient ciphertext $[\![g_{i}^{(t)}]\!]$ as $\cos_{i*}\leftarrow\mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!]),\textrm{for\ }i\in[1,n].$ The smaller $\cos_{i*}$ is, the more likely that $[\![g_{i}^{(t)}]\!]$ is a poisonous gradient.
-Note that $\cos_{i*}\in\mathbb{Z}_{p}$.
-      • $S_{1}$ sets the confidence $co_{i}\leftarrow(p-\cos_{i*})\times cs_{i}$ for the gradient $g_{i}^{(t)},i\in(1,N)$ instead of eliminating outliers, to adapt to non-IID datasets among users.
-      • $S_{1}$ recomputes all users’ credit scores $cs_{i}=cs_{i}\times\lfloor\frac{co_{i}}{co}\rceil,i\in(1,N)$ for the next iteration, where $co\leftarrow\sum_{j=1}^{n}co_{j}$.
-The lower the confidence, the more likely the user is malicious, and thus the user’s credit score needs to be reduced accordingly, and on the contrary, it needs to be increased.
-The credit score contains all history confidences of the user and thus incentives the user to engage in benign training in the whole training process.
-      • $S_{1}$ calculates $co_{i}\leftarrow\lfloor\dfrac{co_{i}}{co}\cdot p\rceil$ and weights the local gradients with their respective confidences and then aggregates them to generate a global gradient $[\![g_{\Sigma}^{(t)}]\!]\leftarrow\sum_{i=1}^{n}[\![co_{i}]\!][\![g_{i}^{(t)}]\!]$, where $g_{\Sigma}^{(t)}=\sum_{i=1}^{n}co_{i}g_{i}^{(t)}$.
-  4. 4.
-    Model Update.
-$S_{1}$ sends $co,[\![g_{\Sigma}^{(t)}]\!]$ and $[g_{\Sigma}^{(t)}]_{s_{i}}\leftarrow\mathbf{FHE.PartDec}_{sk_{s_{i}}}([\![g_{\Sigma}^{(t)}]\!])$ to $U_{i}$.
-Then $U_{i}$ calculates the plaintext aggregated gradient $g_{\Sigma}^{(t)}$ by calling $\mathbf{FHE.PartDec}$ and $\mathbf{FHE.FullDec}$ with $sk_{u_{i}}$.
-Finally, $U_{i}$ updates its local model as $W^{(t+1)}\leftarrow W^{(t)}-co^{-1}\eta^{(t)}g_{\Sigma}^{(t)}$, where $\eta^{(t)}$ is the learning rate.
-TABLE III: Computational costs.
-Schemes Setup Local Training Privacy-Preserving Defense Model Update Total
-$O(\varepsilon+(n+1)\cdot ks$ $O(nlT\cdot train$ $O(4nlT(N\cdot rnd+penc$ $O(3nlT\cdot pdec$ $O(\varepsilon+(n+1)\cdot ks$
-ShieldFL $+l\cdot penc$ $+nlT\cdot penc))$ $+9\cdot pdec+padd$ $+nlT\cdot update)$ $+nlT(6N^{4}\cdot mul+40N^{3}\cdot add$
-$+(2n+1)l\cdot pdec)$ $+N^{2}\cdot pmul))$ $+8N^{3}\cdot rnd+train+update)$
-$O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks$ $O(nlT\cdot train$ $O(n\tau T(3q\cdot rnd+3fenc$ $O(3n\tau T\cdot fdec$ $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks$
-Our scheme $+\tau\cdot fenc$ $+n\tau T\cdot fenc))$ $+9fdec+5fadd$ $+nlT\cdot update)$ $+n\tau T(35n_{f}q\cdot mul+59n_{f}q\cdot add$
+The server $S_{1}$ determines whether the received encrypted local gradients $[\![g_{i}^{(t)}]\!]=\{[\![g_{i,1}^{(t)}]\!],\cdots,[\![g_{i,\tau}^{(t)}]\!]\}$ are regularized by invoking the verification function $\mathbf{ESecJudge}([\![g_{i}^{(t)}]\!])$.
+Only those passing the verification are considered trusted and included in the index set $\mathcal{I}_{\text{trusted}}$.
+  • Poisonous Gradient Baseline Identification.
+For all trusted gradients $\{[\![g_{i}^{(t)}]\!]\}_{i\in\mathcal{I}_{\text{trusted}}}$, the server computes the cosine similarity with the previously aggregated global gradient $[\![g_{\Sigma}^{(t-1)}]\!]$ as $\cos_{i\Sigma}\leftarrow\mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!]),\quad\text{for }i\in\mathcal{I}_{\text{trusted}}.$ The gradient having the lowest similarity score is selected as the poisonous baseline $[\![g_{*}^{(t)}]\!]$.
+  • Byzantine-Tolerant Aggregation.
+To defend against adversarial behaviors while incentivizing honest participants, the server constructs a dual-scoring aggregation strategy using both confidence and credit scores.
+    • Confidence Computation. For each user $i\in\mathcal{I}_{\text{trusted}}$, the cosine similarity between their local gradient and the poisonous baseline is computed as
+      $\cos_{i*}\leftarrow-\ \mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!]).$
+      These similarity values are then normalized to obtain confidence scores:
+      $co_{i}\leftarrow\frac{\exp(\cos_{i*}-\max_{j}\cos_{j*})}{\sum_{j\in\mathcal{I}_{\text{trusted}}}\exp(\cos_{j*}-\max_{k}\cos_{k*})}.$
+    • Credit Score Update. Each participant¡¯s credit score $cs_{i}$ is updated with exponential smoothing
+      $cs_{i}\leftarrow\alpha\cdot cs_{i}+(1-\alpha)\cdot co_{i},$
+      where $\alpha\in[0.7,0.95]$ is a decay coefficient to balance historical consistency and current behavior.
+    • Client Aggregation Weights. The clients’ aggregation weights are computed based on the product of credit and confidence scores
+      $w_{i}\leftarrow\frac{cs_{i}\cdot co_{i}}{\sum_{j\in\mathcal{I}_{\text{trusted}}}cs_{j}\cdot co_{j}}.$
+    • Adaptive Client Filtering.
+To further enhance robustness against adversarial manipulation and reduce the influence of anomalous or overconfident updates, the server performs an optional three-stage client filtering strategy before final aggregation:
+      • (i)
+        Adaptive Weight Mixture. The server first interpolates between uniform weighting and the learned robust weights based on the current round index $t$, using a decaying mixture coefficient:
+        $\lambda\leftarrow\max\left(0,1-\frac{t-T_{\text{warmup}}}{T_{\text{total}}-T_{\text{warmup}}}\right),$
+        $\tilde{w}_{i}\leftarrow\lambda\cdot\frac{1}{n}+(1-\lambda)\cdot w_{i},$
+        where $\tilde{w}_{i}$ denotes the adjusted aggregation weight for client $i$, and $n$ is the total number of clients.
+This allows the server to rely more on uniform weighting in early rounds and gradually transition to confidence-driven weighting as training progresses.
+Clients with $\tilde{w}_{i}<\theta$ (e.g., $\theta=0.05$) are excluded from the aggregation.
+      • (ii)
 […]
-TABLE IV: Communication costs.
-Schemes Setup Local Training Privacy-Preserving Defense Model Update Total
-ShieldFL $O((2n+2)\varepsilon+2nlC_{p})$ $O(nlTC_{p})$ $O(nT(10l+9)C_{p})$ $O(2nlTC_{p})$ $O(2n\varepsilon+13nlTN^{2})$
-Our scheme $O((2n+2)n_{f}\varepsilon^{\prime}+2n\tau C_{f})$ $O(n\tau TC_{f})$ $O(nT(2\tau+4)C_{f})$ $O(2n\tau TC_{f})$ $O(2nn_{f}\varepsilon^{\prime}+5n\tau Tq)$
+      • (iii)
+        Credit Penalty for Filtered Clients.
+To enhance long-term robustness, clients filtered out during either the Normalization Judgment or Outlier Suppression stages are penalized by applying differentiated scaling factors to their credit scores.
+Specifically, a multiplicative decay $\gamma_{1}\in[0,1]$ is applied to clients excluded by normalization, while an inflation factor $\gamma_{2}>1$ is used for clients removed during outlier suppression.
+This adjustment modifies each client’s trust trajectory in future rounds accordingly.
+    Finally, the weighted aggregation of encrypted gradients is computed as
+    $[\![g_{\Sigma}^{(t)}]\!]\leftarrow\sum_{i\in\mathcal{I}_{\text{select}}}\frac{w_{i}}{\sum_{j\in\mathcal{I}_{\text{select}}}w_{j}}\cdot[\![g_{i}^{(t)}]\!].$
+4) Model Update: The server $S_{1}$ partially decrypts and sends the share $[g_{\Sigma}^{(t)}]_{s_{i}}\leftarrow\mathbf{FHE.PartDec}_{sk_{s_{i}}}([\![g_{\Sigma}^{(t)}]\!])$ to each participant $U_{i}$.
+Each user completes decryption using its private key $sk_{u_{i}}$
+$g_{\Sigma}^{(t)}\leftarrow\mathbf{FHE.FullDec}(\{[g_{\Sigma}^{(t)}]_{s_{i}}\},sk_{u_{i}}),$
+and updates its local model using
+$W^{(t+1)}\leftarrow W^{(t)}-\eta^{(t)}\cdot g_{\Sigma}^{(t)},$
+where $\eta^{(t)}$ is the learning rate of round $t$.
+Our proposed privacy-preserving defense is designed to be robust against both conventional data poisoning attacks (as examined in [26]) and more sophisticated adaptive model poisoning attacks, including the AGR-tailored and AGR-agnostic strategies introduced in [42].
+To counter these evolving threats, our defense integrates several key mechanisms that jointly enhance resilience:
+  • Dual-Scoring Trust Assessment.
+A two-tier scoring system is employed, combining per-round cosine-based confidence scores and long-term credit scores.
+This joint quantification reflects both short-term gradient consistency and historical client reliability, effectively preventing adaptive adversaries from persistently imitating benign behaviors to evade detection.
+  • Exponential Credit Score Smoothing.
+Credit scores are updated via exponential smoothing with decay factor $\alpha\in[0.7,0.95]$, enabling the system to gradually penalize clients with recurrent suspicious updates while remaining tolerant to occasional noise in benign gradients.
+  • Soft Confidence Normalization.
+Instead of applying hard thresholding, we utilize softmax-based normalization of negative cosine similarities to compute smooth confidence scores.
+This soft treatment maintains sensitivity to fine-grained deviations and provides robustness against gradient-level perturbations crafted by adaptive attackers.
+  • Adaptive Weight Mixing and Thresholding.
+An adaptive mixing scheme transitions from uniform weighting to fully trust-weighted aggregation as training progresses, enhancing early-round stability.
+Clients with extremely low trust weights (below threshold $\theta$) are excluded from aggregation to reduce stealth attack surfaces.
+  • Outlier Suppression via Drop Detection.
+Our outlier suppression step explicitly targets highly confident yet anomalously extreme updates, which are likely crafted to exploit the statistical patterns of benign gradients for malicious purposes.
+This mechanism is particularly critical for defending against adaptive adversaries, such as those employing AGR-tailored strategies.
+In such attacks, malicious gradients are deliberately constructed based on the distributional characteristics of benign updates, typically by adding a scaled deviation vector to their mean.
+These adversarial updates often evade baseline filters due to their close proximity to the benign gradient space.
+Furthermore, they may achieve abnormally high confidence scores by aligning even more closely with the directionality of benign updates than the benign clients themselves.
+  • Credit Penalty for Filtered Clients.
+This mechanism addresses the risk of adaptive adversaries regaining trust too easily after being filtered.
+Specifically, clients removed during the Normalization Judgment step are presumed to have produced non-normalized or low-quality gradients and thus receive a credit decay to reduce their future influence.
+In contrast, clients filtered via the Outlier Suppression mechanism may have presented gradients that appear highly confident yet abnormally extreme.
+These gradients often arise from adversarial strategies that construct malicious updates by scaling deviations from the benign gradient mean, making them closely aligned with benign directionality and capable of evading baseline checks.
+Since such clients may initially appear trustworthy, a reverse penalty (i.e., increasing their credit score using $\gamma_{2}>1$) makes them more likely to be flagged again in subsequent outlier filtering, reinforcing detection consistency across rounds.
 6 Theoretical Analysis
@@
 […]
-6.2 Performance Comparison
-Our scheme has lower computational and communication burdens compared with ShieldFL.
-We compare the complexity of the two schemes in their split processes and the comparison results are shown in Table III and Table IV.
-TABLE V: Computational costs of the encryption algorithm in ShieldFL and our scheme.
-Schemes Encryption Decryption Homo. Addition Homo. Multiplication
-ShieldFL $penc$ $pdec$ $padd$ $pmul$
-$O(N(mN\cdot mul+add))$ $O(N((N+1)\cdot mul+add))$ $O(N^{2}\cdot mul)$ $O(2N\cdot rnd+(4mN^{2}+9N^{2}+6N)mul+10N\cdot add)$
-Our scheme $fenc$ $fdec$ $fadd$ $fmul$
-$O(n_{f}q\cdot(2mul+3add))$ $O(n_{f}q\cdot(mul+3add))$ $O(n_{f}q\cdot add)$ $O(n_{f}q\cdot(5mul+2add))$
-TABLE VI: Communication costs of the encryption algorithm in ShieldFL and our scheme
-Schemes Homo. multiplication Ciphertext size
-ShieldFL $C_{pmul}:O(5N^{2})$ $C_{p}:O(N^{2})$
-Our scheme 00 $C_{f}:O(p)$
-  • In Setup process of both schemes, KC and $S_{1}$ distribute keys and initial model, respectively.
-The computation complexities of ShieldFL and our scheme are $O(\varepsilon+(n+1)\cdot ks+l\cdot penc+(2n+1)l\cdot pdec)$ and $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks+\tau\cdot fenc+(2n+1)\tau\cdot fdec)$, respectively, where $ks$, and $penc/fdec$, and $pdec/fdec$ represent the running time when implementing $\mathbf{PHE.KeySplit},$ $\mathbf{PHE.Enc}/$$\mathbf{FHE.Enc},$ and $\mathbf{PHE.Dec}/$$\mathbf{FHE.Dec}$, respectively.
-The communication complexities for ShieldFL and our scheme are $O((2n+2)\varepsilon+2nlC_{p})$ and $O((2n+2)n_{f}\varepsilon^{\prime}+2n\tau C_{f})$, respectively, where $C_{p}$/$C_{f}$ represents the ciphertext sizes of the two-trapdoor PHE/FHE.
-  • In Local Training process of both schemes, $n$ users train their models locally, encrypt their local gradients, and transmit them to the server.
-The computation complexities of ShieldFL and our scheme are $O(nlT\cdot train+nlT\cdot penc))$ and $O(nlT\cdot train+n\tau T\cdot fenc))$, respectively, where $T$ is the total training iterations.
-The communication complexities of ShieldFL and our scheme are $O(nlTC_{p})$ and $O(n\tau TC_{f})$, respectively.
-  • In Privacy-Preserving Defense process, the computation and communication costs of both schemes mainly come from the protocols $\mathbf{SecJudge}/\mathbf{ESecJudge}$ and $\mathbf{SecCos}/\mathbf{ESecCos}$.
-For ShieldFL, the computation and communication complexities for executing the two protocols are $O(4nlT(N\cdot rnd+penc+9pdec+padd+N^{2}\cdot pmul))$ and $O(nT(10l+9)C_{p})$, respectively, where $rnd,padd$ and $pmul$ represent the running time of generating a random noise, a homomorphic addition, and a homomorphic multiplication, respectively.
-For the enhanced scheme, the computation and communication complexities for executing the two protocols are $O(n\tau T(3q\cdot rnd+3fenc+9fdec+5fadd+3fmul))$ and $O(nT(2\tau+4)C_{f})$, respectively, where $fadd/fmul$ represent the running time of a fully homomorphic addition/multiplication.
-  • In Model Update process, all users interact with $S_{1}$ to obtain the decrypted aggregated gradient and then update their local models.
-For ShieldFL, the computation and communication complexities of this process are $O(nlT(3\cdot pdec+update))$ and $O(2nlTC_{p})$, respectively.
-For our scheme, the computation and communication complexities are $O(3n\tau T\cdot fdec+nlT\cdot update)$ and $O(2n\tau TC_{f})$, respectively.
-The total computational costs of ShieldFL and our scheme are $O(\varepsilon+(n+1)\cdot ks+nlT(8penc+36pdec+4padd+4N^{2}\cdot pmul+4N\cdot rnd+train+update)$ and $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks+n\tau T(7fenc+9fdec+5fadd+3fmul+3q\cdot rnd)+nlT\cdot(train+update))$, respectively.
-Table V presents the computational costs of $penc,pdec,padd,pmul,fenc,fdec,fadd,fmul$ related to general addition $add$ and general multiplication $mul$ operations.
-Thus, the total computational costs of ShieldFL and our scheme are $O(\varepsilon+(n+1)\cdot ks+nlT(36N^{4}\cdot mul+40N^{3}\cdot add+8N^{3}\cdot rnd+train)$ and $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks+n\tau T(35n_{f}q\cdot mul+59n_{f}q\cdot add+9q\cdot rnd)+nlT\cdot train)$, respectively.
-The total communication costs of ShieldFL and the enhanced scheme are $O(2n\varepsilon+13nlTC_{p})$ and $O(2nn_{f}\varepsilon^{\prime}+5n\tau TC_{f})$, respectively.
-Table VI presents the communication costs of homomorphic multiplication operations and the ciphertext sizes of the two schemes.
-Our scheme’s homomorphic multiplication operation is performed locally, and thus its communication cost is 00.
-The homomorphic multiplication protocol in ShieldFL requires interactions between the two servers, during which five ciphertexts are transformed.
-Thus, its communication cost is $O(5N^{2})$.
-Therefore, the total communication costs of ShieldFL and the enhanced scheme are $O(2n\varepsilon+13nlTN^{2})$ and $O(2nn_{f}\varepsilon^{\prime}+5n\tau Tq)$, respectively.
+6.2 Convergence Analysis
+We provide a detailed theoretical convergence analysis of the proposed PBFL scheme in Appendix C.
+Specifically, we decompose the aggregation error into three components—(i) residual Byzantine bias, (ii) stochastic gradient noise, and (iii) gradient diversity—arising respectively from malicious updates, local sampling noise, and data heterogeneity.
+We then formally show that, under standard smoothness and bounded-variance assumptions, the proposed algorithm achieves the same sub-linear convergence rate $\mathcal{O}(1/\sqrt{T})$ as classical FedAvg, up to an additive error term that is proportional to the Byzantine fraction $\beta$ and decreases as the number of trusted participants increases.
+This result theoretically confirms the robustness and efficiency of our defense mechanism.
+6.3 Performance Analysis
+We provide a detailed performance comparison of computational and communication complexities between our proposed scheme and ShieldFL, which is presented in Appendix D.
+The analysis covers all major phases including Setup, Local Training, Privacy-Preserving Defense, and Model Update.
+Our results show that, due to ciphertext computation simplicity and cryptographic primitive efficiency, our scheme significantly outperforms ShieldFL in both computation and communication efficiency.
+Specifically:
+  • Computational Complexity: Our scheme reduces asymptotic encryption, decryption, and homomorphic operation costs by a factor related to the smaller ciphertext dimension ($n_{f}q\ll N^{2}$).
+This translates to a lower overhead in all training rounds.
+  • Communication Cost: Since homomorphic multiplication in our secure computation methods $\mathbf{ESecCos}$ and $\mathbf{ESecJudge}$ is executed locally without inter-party interaction, its communication cost is effectively zero.
+In contrast, ShieldFL incurs significant overhead due to interactive homomorphic multiplication in their secure computation protocols.
+6.4 Discussion and Practical Considerations
+While our enhanced scheme achieves significant improvements in privacy and efficiency, several practical considerations remain for real-world FL deployments.
+This section discusses two key aspects: the feasibility of decentralizing the Key Center and the challenges of defending against adaptive and malicious adversaries.
+(1) Decentralizing the Key Center: Our current design assumes a trusted centralized Key Center for key management and distribution.
+This assumption, however, may not always hold in heterogeneous or cross-silo FL environments where participants belong to different administrative domains with varying trust levels.
+To mitigate this limitation, decentralized approaches such as threshold cryptography and Distributed Key Generation (DKG) protocols can be adopted.
+These techniques distribute key management responsibilities among multiple independent parties, thereby eliminating the single point of trust and reducing the risk of key compromise.
+In addition, blockchain-based solutions and secure multiparty computation (MPC) protocols could be integrated to provide transparent, auditable, and tamper-resistant key management.
+While these decentralized mechanisms can enhance robustness and trustworthiness, they inevitably introduce challenges such as increased communication overhead and coordination complexity, which require further study.
+(2) Defending Against Adaptive and Malicious Threats: Our security analysis primarily considers malicious clients and at most one corrupted (malicious) server, ensuring the privacy of honest clients’ data under the assumption that the two servers do not collude.
+However, in real-world FL deployments, adversaries may behave both maliciously and adaptively, dynamically deviating from the protocol to launch sophisticated attacks such as adaptive model poisoning or adaptive inference.
+We explicitly evaluate our framework under an adaptive model poisoning threat, where the adversary dynamically adjusts its malicious gradients based on the gradients of benign clients before injecting them into the training process.
+This adaptive strategy makes detection substantially more difficult.
+Our experimental results demonstrate that the proposed framework can effectively mitigate this class of adaptive attacks.
+Beyond model poisoning, more advanced adaptive threats have been reported, including adaptive backdoor attacks that iteratively modify backdoor triggers to evade defenses and adaptive inference attacks that progressively extract private information from model updates.
+These attacks evolve over training rounds and present challenges that our current design does not fully address.
+Future work will focus on strengthening the framework against these stronger adversarial models by extending robustness beyond poisoning to cover adaptive backdoor and inference-based attacks.
+Potential directions include developing adaptive anomaly detection frameworks that dynamically adjust detection thresholds or defense strategies based on adversarial behaviors observed during training (e.g., online learning-based detectors that evolve as attacks change) and incorporating temporal consistency checks and cross-round validation mechanisms to detect adversaries that gradually adapt their malicious updates to evade static defenses.
+Furthermore, leveraging ensemble or diversity-based defenses (where multiple defense mechanisms operate in parallel and dynamically reconfigure based on the detected threat level) could make it substantially harder for adaptive attackers to circumvent all defenses simultaneously.
+Beyond adaptive attacks, we also plan to address the threat of malicious servers that compromise computation correctness by integrating verifiable computation or zero-knowledge proofs to guarantee both privacy and correctness even when a server is malicious.
 7 Conclusion